### v1

# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

Abstract
Recent advancements in large language models (LLMs) have facilitated a wide range of applications with distinct quality-of-experience requirements, from latency-sensitive online tasks, such as interactive chatbots, to throughput-focused offline tasks like document summarization. While deploying dedicated machines for these services ensures high-quality performance, it often results in resource underutilization.
This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements. HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor for batch execution time estimation and an SLO-aware profiler to quantify interference, and (2) SLO-aware offline scheduling policies that maximize throughput and prevent starvation, without compromising online serving latency. Our evaluation on production workloads shows that HyGen achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency. The code will be released soon.
1 Introduction
Large language models (LLMs) have emerged as transformative tools across diverse domains, powering applications from interactive chatbots to document analysis systems Grand View Research (2023). As LLM services continue to spread, their serving deployment must handle both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al. (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)). Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), to ensure smooth user interactions. In contrast, offline tasks prioritize high throughput and resource utilization, often processing large batches of requests with relaxed latency constraints. The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al. (2023); Zhao et al. (2024); Patel et al. (2024); Stojkovic et al. (2024).
However, real-world LLM workloads demonstrate significant temporal variations in request loads. Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3). To maintain acceptable latency under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster. By opportunistically co-locating offline tasks with online tasks during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests. However, realizing this opportunity requires addressing several fundamental challenges.
First, LLM services exhibit diverse latency requirements across applications Liu et al. (2024a). Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency. These requirements often manifest in different statistical metrics—from strict P99 latency bounds to mean performance targets—making it difficult to establish unified resource-sharing policies.
Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands. Online requests arrive in unexpected bursts, while offline tasks appear in irregular batch intervals. This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
Third, co-locating online and offline workloads introduces interference, which degrades the performance of both. For instance, offline requests can delay time-sensitive online requests, while using large batch sizes to boost offline request throughput can negatively impact the latency of online requests. Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads. HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of heterogeneous request batches, (2) an interference-aware profiler that quantifies the performance impact of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online serving.
Our evaluation on production workloads shows that HyGen guarantees latency requirements in co-locating requests and achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency.
To summarize, this paper makes the following contributions:
  • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with guaranteed performance isolation.
  • To improve the performance isolation and throughput of HyGen, we design performance control mechanisms and SLO-aware offline scheduling policies.
  • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
2 Related Work
LLM Online Serving.
Recent works have improved GPU computation through optimized kernels Dao et al. (2022); Ye et al. (2025), memory efficiency through efficient KV-cache management techniques Lin et al. (2024a); Liu et al. (2024b); Qin et al. (2024), distributed inference through various model parallelism strategies Shoeybi et al. (2019); Pope et al. (2023); Liu et al. (2023), and model compression through quantization approaches Xiao et al. (2023); Lin et al. (2024b); Dettmers et al. (2022). Orca Yu et al. (2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching, and enables early exits for completed requests, thereby reducing batching latency. vLLM Kwon et al. (2023) develops PagedAttention for efficient memory management, while Sarathi-Serve proposes chunked prefill to minimize inter-request interference within batches. However, few existing systems explicitly address the challenge of co-locating latency-critical online requests with offline workloads to maximize resource utilization—a key focus of our work.
Throughput-Oriented LLM Inference.
As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al. (2022) and FlexGen Sheng et al. (2023) enable efficient LLM serving on commodity GPUs by offloading strategies. FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput. For KV cache optimization, InfiniGen Lee et al. (2024) proposes dynamic cache management through selective prefetching based on token importance prediction. S3 Jin et al. (2023) tackles the KV cache problem through output sequence length prediction, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length. NanoFlow Zhu et al. (2024) focuses on maximizing hardware utilization through intra-device parallelism, introducing nano-batching and execution unit scheduling to overlap compute-, memory-, and network-intensive operations.
Co-locating Workloads.
Co-locating latency-critical applications with batch applications is a widely adopted approach to improve resource utilization in datacenters, with the popular resource partitioning techniques Chen et al. (2019b); Zhu & Erez (2016); Lo et al. (2015); Wu & Martonosi (2008); Chen et al. (2019a); Kasture et al. (2015); Sanchez & Kozyrakis (2011); Kasture & Sanchez (2014); Zhang et al. (2019); Iorgulescu et al. (2018); Patel & Tiwari (2020); Chen et al. (2023). Prior work has also explored multiplexing to serve multiple DNNs Dhakal et al. (2020); Tan et al. (2021); Zhang et al. (2023b), LLMs Duan et al. (2024), or LoRAs Chen et al. (2024); Sheng et al. (2024a); Wu et al. (2024). However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
3 Background and Motivation
This section first introduces the core concepts of LLM serving systems (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
3.1 LLM Serving
LLM generation consists of two sequential stages: (i) Prefilling Stage: It processes the entire input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) of the input sequence to avoid regeneration in subsequent tokens, and then the (ii) Decoding Stage: the model generates tokens one at a time in an auto-regressive manner. Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead. This stage iterates until meeting a termination condition, such as generating an end-of-sequence token or reaching the maximum length.
Model executions in the prefilling stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously. Yet in the decoding phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al. (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
Serving Workloads.
LLM serving deployment can be categorized into two primary scenarios. The first scenario is online serving, which targets real-time user interactions, such as chatbots, code assistants, and interactive applications OpenAI (2022); Microsoft (2023). This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
The second counterpart is offline serving, which prioritizes throughput over latency, measured in requests per second (RPS) or tokens per second (TPS) for prefill/decode operations. Key applications include model benchmarking Hendrycks et al. (2021b); Meta-Team (2024), document processing Chen et al. (2021); Zhang et al. (2023a), data cleaning Narayan et al. (2022); Zheng et al. (2023), and synthetic data generation Allal et al. (2024).
3.2 Motivation and Challenges
(a) 1 hour
(b) 2 minutes
Figure 1: Request rate variation in Microsoft Azure’s LLM service over one-hour and two-minute periods, showing significant temporal fluctuations between peak and off-peak periods.
LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads. Figure 1 reports our analysis of Microsoft Azure’s LLM service Patel et al. (2024), showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns. This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
This load variability suggests an opportunity to improve GPU utilization by co-locating online inference requests with offline batch workloads. During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources. However, co-locating these workloads introduces several fundamental challenges:
  1. 1.
    Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements. For example, users of different ages have different reading speeds, thus TBT requirements Liu et al. (2024a). These requirements are often expressed through multiple statistical metrics - some services prioritize tail latency guarantees (e.g., P99 TTFT $\leq$ 1000ms), others focus on average performance (e.g., mean TBT $\leq$ 50ms), and many need to satisfy both simultaneously. How to provide latency guarantees for diverse requests while sharing resources?
  2. 2.
    Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths. A seemingly simple query might trigger an elaborate response requiring thousands of tokens. How to perform efficient scheduling in the wild?
  3. 3.
    Request Interference: Co-locating online and offline workloads introduces interference that degrades system performance. Large batches of offline requests arriving simultaneously can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests. Worse, batching requests of long inputs (e.g., a long document summarization request with 2,048 tokens) with short interactive queries (e.g., 128 tokens) can result in a dramatic increase in attention computation overhead, elongating the latency for all requests in the batch by an order of magnitude Agrawal et al. (2024). How to account for interference in co-locating requests?
4 The HyGen Design
4.1 Overview
Figure 2: HyGen overview: Online requests and offline requests sent by users are asynchronously selected and processed.
In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees. Unlike conventional LLM serving systems that segregate online and offline workloads onto separate GPU clusters, HyGen coordinates both types of requests through a two-phase scheduling approach, maximizing resource utilization during off-peak periods while preserving performance isolation.
Algorithm 1 HyGen SLO-aware scheduler
1:  function SLO_AWARE_SCHEDULE
2:     Input: running requests $R$, request queue $Q$,
3:      latency budget $t$, remaining chunk size $c$,
4:      memory budget $m$
5:     Output: batched requests $B$
6:     $B\leftarrow\{\}$
7:     for $r\in R$.decoding do
8:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
9:        if $t_{req}\leq t$ then
10:           $t\leftarrow t-t_{req}$
11:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        end if
13:     end for
14:     for $r\in R$.prefilling $\cup Q$ do
15:        TRY_SCHEDULE:
16:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
17:        if $l>0$ then
18:           $t\leftarrow t-t_{req}$
19:           $c\leftarrow c-l$
20:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
21:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
22:        else
23:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
24:              PERFORM_PREEMPTION($R$, $m$)
25:              goto TRY_SCHEDULE
26:           else
27:              break
28:           end if
29:        end if
30:     end for
31:     return $B$
32:  end function
Figure 2 illustrates the system architecture of HyGen. To handle diverse latency requirements and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector that tracks active requests and their states. This separation allows HyGen to support scheduling policies for different request types without the complexity of managing mixed priorities in a single queue.
The scheduling process in HyGen operates in two consecutive phases to carefully manage interference between online and offline requests. In the online phase, the scheduler processes the online request queue to construct an initial batch based on latency SLOs. This phase follows traditional continuous batching principles and supports various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al. (2023) and Virtual Token Counter (VTC) Sheng et al. (2024b). The scheduler prioritizes running requests before processing the waiting queue, continuing until reaching memory, batch size, or chunk size limits. A detailed workflow of this process is displayed in Algorithm 1. During a default two-phase scheduling, this function is called two times separately for online and offline requests.
During online batch establishment, when the current online batch reaches the memory limit and more online requests remain to be scheduled, HyGen performs a priority-based preemption to maximize SLOs. To handle dynamic request arrivals while maintaining service guarantees, we implement an online-first preemption mechanism, where online requests are prioritized over all offline requests to minimize interference. Offline requests are prioritized based on their original priority in the waiting queue and their memory consumption. Preempted offline requests are removed from the current running list and reinserted into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
After establishing the online batch, the scheduler enters the offline phase where it opportunistically fills the remaining capacity with offline requests. By utilizing a latency predictor to estimate batch execution time and determine remaining latency budget, instead of relying on static chunk size or batch size limits, we achieve both higher accuracy in interference prediction and strong transferability across different hardware configurations and workload patterns. This enables higher throughput while maintaining SLO guarantees by eliminating the need for conservative scheduling. The selection of offline requests then considers this available batch execution latency budget derived from user-specified SLOs, opportunities for efficient execution through prefix sharing, and current batch composition and resource constraints. Within this phase, running requests are prioritized over waiting ones to maintain execution efficiency.
HyGen addresses these complex scheduling decisions through two key components detailed in the following sections. Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation decisions. Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining performance isolation through careful workload co-location. A detailed algorithm for the entire HyGen scheduling workflow can be found in Appendix A.1.
4.2 Performance Control Mechanisms
Efficient co-location of online and offline requests requires precise control over execution timing and resource allocation. The key challenge lies in accurately estimating the performance impact of scheduling decisions while ensuring end-to-end SLO compliance. This section presents two core mechanisms in HyGen that address this challenge: a latency predictor for estimating batch execution times and an SLO-aware profiler for translating these estimates into scheduling decisions.
Latency Predictor.
The design of our latency predictor is guided by three key requirements. First, it must provide fast inference to support real-time scheduling decisions when evaluating different batch compositions. Second, it needs to be robust across varying workload patterns conditions to maintain reliable performance. Third, it should be adaptable to different hardware configurations to capture their unique performance characteristics.
The execution time of an LLM serving batch is primarily influenced by two distinct processing stages with different computational patterns. The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length. This stage’s total load depends on both the number of requests and their individual sequence lengths. In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens.
To capture these characteristics, we model batch execution time as:
$T_{batch}=f(S_{p},S_{d},S_{p}^{2},S_{d}^{2},N_{p},N_{d})$ (1)
where $S_{p}$ and $S_{d}$ represent the total number of prefill and decode tokens in the batch, respectively. The quadratic terms ($S_{p}^{2}$ and $S_{d}^{2}$) account for non-linear scaling effects, particularly in the prefill phase. $N_{p}$ and $N_{d}$ represent the number of prefill and decode requests in the batch, respectively.
We implement this model using linear regression, chosen for its combination of efficiency and effectiveness. The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions. The training process involves three steps:
  1. 1.
    Training Data Collection: We systematically profile target hardware across diverse batch compositions, varying the number of requests in different phases, sequence length distributions, and total batch sizes.
  2. 2.
    Feature Engineering: For each batch, we compute six key features ($S_{p}$, $S_{d}$, $S_{p}^{2}$, $S_{d}^{2}$, $N_{p}$, $N_{d}$) based on token counts and batch sizes, capturing both linear and quadratic scaling.
  3. 3.
    Model Training: Linear regression coefficients are optimized to minimize prediction error across the training set, providing sufficient accuracy for scheduling decisions while maintaining fast inference times.
SLO-Aware Profiling.
While accurate latency prediction is essential, translating these predictions into scheduling decisions requires careful consideration of end-to-end performance requirements. Our SLO-aware profiler addresses this by determining appropriate latency budgets that ensure SLO compliance under various operating conditions.
The profiler first analyzes specific workload-SLO combinations to establish viable latency budget ranges. Then, given that SLO has a positive correlation with batch latency, we perform the binary search on the batch latency budget to get an upper limit that meets the SLO. This profiling enables three key capabilities. First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types. Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements. Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
Together, these mechanisms enable HyGen to make precise scheduling decisions that maximize resource utilization while maintaining strict performance guarantees. The latency predictor provides the quantitative basis for evaluating scheduling options, while the SLO-aware profiler ensures these decisions translate into consistent end-to-end performance.
4.3 SLO-Aware Offline Scheduling Policies
Offline scheduling in HyGen differs from online scheduling due to the variability of resource availability. Resources allocated to offline requests depend on the residual capacity after scheduling online requests, which fluctuates with the intensity of the online workload. Consequently, offline scheduling policies must dynamically adapt to available resources and optimize their specific objectives under such constraints. Common objectives include maximizing throughput, supporting multi-priority workloads, and, optionally, ensuring fairness Sheng et al. (2024b). The prioritization of these objectives varies depending on the chosen policy.
Prefix sharing Gim et al. (2024); Gao et al. (2024); Zheng et al. (2024a) has emerged as a widely-adopted technique to significantly improve throughput and reduce latency in LLM serving systems. The benefits are particularly pronounced in offline LLM serving, where requests can be strategically ordered to maximize prefix cache reuse Zhao et al. (2024); Zheng et al. (2024b). This optimization is especially effective for offline workloads since they have relaxed latency constraints, allowing the system to accumulate and reorder requests to maximize sharing opportunities. To showcase the extensibility and flexibility of HyGen’s design, we next introduce the SLO-aware Prefix Sharing Maximization (PSM) strategy, a throughput-optimized offline scheduling policy that leverages prefix reuse to improve resource efficiency. Recognizing the potential starvation issue inherent in naive PSM, we extend it with a fairness-enhanced variant. This improved PSM algorithm incorporates mechanisms such as freshness-aware scheduling and a unified utility function to balance throughput and fairness, mitigating starvation without compromising efficiency.
The modular architecture of HyGen enables seamless integration of advanced offline scheduling policies. In addition to PSM, the system supports other strategies, such as a default First-Come-First-Serve (FCFS) policy and a Deadline-First policy for basic scheduling needs. This flexibility allows HyGen to cater to a wide range of objectives, including multi-priority support and fairness, making it a robust and adaptable solution for hybrid serving scenarios with dynamic and heterogeneous workloads.
Prefix Sharing Maximization Strategy.
To fully leverage prefix sharing among offline requests, HyGen employs a Prefix Sharing Maximization (PSM) strategy designed for SLO-aware offline scheduling. This strategy organizes offline requests into a prefix tree and performs Depth-First Search (DFS) to generate a sorted waiting list. HyGen then performs SLO-aware scheduling on offline requests using this sorted list in a similar manner as introduced in Section 4.1, while maintaining the original DFS order to maximize prefix sharing. This enables reducing redundant computation and improving throughput.
Despite its efficiency, the naive PSM strategy can lead to starvation for requests with minimal prefix-sharing potential. This issue can be mitigated by combining requests selected by PSM with those chosen based on freshness. The trade-off between these two selection methods can be managed using a utility ratio, ensuring a balance between efficiency and fairness. During scheduling, offline waiting requests are stored in both a prefix tree for maximum prefix sharing, and a self-balanced binary search tree (BST) sorted by freshness. When a new offline request can be scheduled from the waiting list, an offline request is selected either from the head of the queue in the DFS order of the prefix tree, or from the most stale request in the self-balanced BST, based on the utility value. The other data structure then syncs up by deleting the selected request from their lists, namely by prefix and index lookup. The formalized algorithm is detailed in Appendix A.2.
While starvation due to the strict prioritization of online requests over offline ones is unavoidable to ensure performance isolation, the enhanced PSM algorithm effectively mitigates starvation caused by prefix-sharing optimizations. These enhancements make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
5 Performance Evaluation
Figure 3: HyGen respects latency requirements in co-locating requests.
Figure 4: HyGen improves serving throughput under the same SLOs.
5.1 Evaluation Setup
Implementation.
We implement HyGen on top of the open-source implementation of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code.
Testbeds.
We evaluate HyGen on two server configurations. The first server is equipped with 4 NVIDIA A100 GPUs (40GB VRAM each), while the second server contains 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers are identically configured with an AMD EPYC Milan CPU (1 socket, 64 cores, SMT disabled), 256GB DDR4 RAM, and a 1.5TB NVMe SSD (PCIe 4.0).
Models.
We used Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models to test the performance of HyGen on varying model structures and sizes. For the two models, we used A100 and A40 GPUs, respectively.
Workloads.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour trace taken from the LLM services in Azure with real-world requests and timestamps. The request rate variation is shown in Figure 1. We randomly sampled the original trace to get the desired overall QPS that suits our hardware serving capacity. Specifically, for a time duration of $T$ seconds, we would sample $T\times Q$ requests from this time window to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for summarization of long documents.
Metrics.
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. For throughput evaluation, we evaluated our system on the maximum throughput of offline requests while meeting interference SLOs.
Baselines.
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first two-phase scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to be an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, namely mean TBT, P99 TBT, mean TTFT, and mean TBT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
Figure 5: Prefix sharing maximum scheduler achieves higher serving capacity.
Figure 6: SLO-aware profiler bridges the gap between batch latency and overall QoE.
Figure 7: HyGen latency predictor achieves high accuracy for batch latency prediction.
Figure 8: HyGen meets stringent SLO under various QPS settings
Figure 9: Hygen dynamically controls offline throughput according to online workload and latency budget.
(a) SLOs
(b) Throughput
Figure 10: Hygen meets SLOs and achieves higher throughput for Yi-34B model using TP=2, PP=2.
(a) SLOs
(b) Throughput
Figure 11: Hygen meets SLOs and achieves higher throughput for Mooncake trace and Mistral-7B model.
(a) SLOs
(b) Throughput
Figure 12: Hygen meets SLOs and achieves higher throughput for CNN/DM offline trace.
Figure 13: HyGen is able to meet multiple SLOs.
Figure 14: HyGen design is robust to predictor accuracy.
Figure 15: Hygen adjusts offline throughput based on online QPS.
HyGen improves serving throughput. Figure 4 showcases the offline throughput of HyGen for various metrics and tolerance ratios. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.37×5.37$\times$5.37 × to 5.84×5.84$\times$5.84 × throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler achieve higher offline serving efficiency compared to their simplified counterparts.
5.3 Performance Breakdown
Accuracy of latency predictor.
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 7 shows the CDF of absolute error and absolute percentage error of the latency predictor. HyGen latency predictor is able to achieve a mean absolute percentage error of only 1.78% and 1.07%, verifying the high accuracy of our predictor design.
Impact of prefix sharing.
To test HyGen’s compatibility with prefix sharing, we did a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset, running on Llama2-7B model. In our simulation, the shared prompt prefix length between two consecutively scheduled offline requests is deducted from the original prompt length of the latter one to simulate prefix sharing. We tested the offline throughput of HyGen with and without our prefix-aware offline schedule policy. Figure 6 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix-aware scheduling policy.
Impact of SLO-aware profiler.
To demonstrate the effect of the SLO-aware profiler in HyGen, we conducted experiments where we simply set the desired mean TBT SLO as the batch latency budget. Figure 6 shows the performance gap between individual batch latencies and overall mean TBT, and that our SLO-aware profiler is able to bridge this gap and provide hybrid serving with controlled SLO.
Breakdown by time.
Figure 9 shows a temporal breakdown of the throughput performance of HyGen during one serving process. At runtime, HyGen shows the ability to dynamically adjust offline throughput according to current online workload and overall latency budget. HyGen batches offline requests more aggressively during troughs of online QPS and lowers offline throughput during online bursts, harnessing compute resources in an adaptive manner.
5.4 Ablation Studies
Impact of model parallelisms.
To evaluate HyGen’s effectiveness in distributed inference scenarios, we deployed the Yi-34B model Young et al. (2024) using a hybrid parallelization strategy: tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. The online and offline workloads are Azure LLM Inference trace and arXiv summarization, respectively. Figure 10(a) and Figure 10 show that HyGen maintains its ability to meet SLOs and achieves higher throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
Impact of SLO requirements.
To further evaluate HyGen’s ability to meet stringent SLOs under various settings, we conducted experiments on the four aforementioned metrics under varied QPS settings, each with a 5% interference tolerance ratio. Figure 8 shows that HyGen is able to meet stringent SLOs for all metrics under various QPS settings.
We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. Here, we tested HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%. Figure 13 shows that at a lower mean TBT tolerance, the performance of HyGen is bounded by mean TBT SLOs; however, after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control, demonstrating the flexibility of HyGen in meeting multiple SLOs.
Impact of models and datasets.
We further evaluate HyGen on Mistral-7B and MooncakeQin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace. The offline dataset used in the first experiment is arXiv summarization. For the second experiment, we replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015) and used Llama2-7B and Azure LLM Inference trace as the model and online dataset, respectively. Figure 11 and Figure 12 show HyGen achieves superior performance than its counterparts under these two settings.
Impact of predictor accuracy.
To further evaluate the robustness of HyGen, we tested HyGen using several pre-trained LR latency predictors taken from other settings and workloads and tested them on Azure LLM Inference trace, resulting in varied prediction accuracy. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. We notice that HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on our testing CPU, guaranteeing transferability and negligible overhead.
Impact of online arrival rates.
Under high online request arrival rates, online requests may approach system serving capacity, leaving little headroom for offline request co-location. Figure 15 shows how online QPS affects offline throughput under a 5% P99 TBT SLO tolerance ratio. As online load increases, HyGen is able to adjust offline throughput according to the residual serving capacity of the system, while maintaining higher throughputs.
6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ a control mechanism to predict and manage interference impacts, while the scheduling framework supports multiple optimization objectives for offline serving. Evaluation on production workloads demonstrates that HyGen achieves up to 5.84$\times$ higher throughput while maintaining strict latency SLOs.
Appendix
Appendix A Algorithms
A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed explanation of the two-phase scheduling algorithm introduced in Section 4. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the coming scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization size $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
Algorithm 2 HyGen two-phase scheduler
1:  global $Q_{on}$ (online request queue), $Q_{off}$ (offline request queue)
2:  global $R_{on}$ (online request running list), $R_{off}$ (offline request running list)
3:  global $Q_{send}$, $Q_{recv}$ (message queue)
4:  global $L$ (latency budget), $M$ (memory budget), $C$ (chunk size), $M_{off}$ (offline memory)
5:  function ASYNC_SCHEDULER
6:     while True do
7:        scheduled requests $S\leftarrow Q_{send}$.get(block=True)
8:        UPDATE_REQUEST_STATUS($S$, $R_{on}$, $R_{off}$)
9:        batched requests $B\leftarrow$}
10:        latency budget $t\leftarrow L$
11:        memory budget $m\leftarrow$ GET_FREE_MEMORY() + $M_{off}$
12:        chunk size $c\leftarrow C$
13:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{on}$, $Q_{on}$, $t$, $c$, $m$)
14:        if $m<M_{off}$ then
15:           PREEMPT_OFFLINE($R_{off}$, $m$, $M_{off}$)
16:        end if
17:        $m\leftarrow m-M_{off}$
18:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
19:        $Q_{recv}$.send($B$, $t$, $m$)
20:     end while
21:  end function
22:  function SCHEDULER
23:     Output: batched requests $B$
24:     scheduled requests $B$, latency budget $t$, memory budget $m$leftarrow $Q_{recv}$.get(block=True)
25:     if REQUEST_ARRIVAL then
26:        UPDATE_BUDGET($B$, $t$, $m$)
27:        if $t<0$ or $m<0$ then
28:           PREEMPT_UNTIL_FIT($B$, $t$, $m$)
29:        end if
30:     end if
31:     $Q_{send}$.send($B$)
32:     return $B$
33:  end function
A.2 SLO-Aware Prefix Sharing Maximization Algorithm
This section details the SLO-aware prefix sharing algorithm design in HyGen. For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness. A utility value between 0 and 1 is used to balance the chance between these two data structures. During scheduling, prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, and keeps their original order when scheduling running requests, effectively utilizing prefix sharing while avoiding starvation.
Algorithm 3 Prefix-sharing-aware offline scheduler
1:  Construct prefix tree $T_{p}$ and self-balanced BST $T_{f}$
2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$, utility value $u$
4:     Output: batched requests $B$
5:     for $r\in R$ do
6:        if $r$.state == DECODE then
7:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
8:           if $t>t_{req}$ then
9:              break
10:           end if
11:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
12:        else if $r$.state == PREFILL then
13:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
14:           if $l>0$ then
15:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
16:           else
17:              break
18:           end if
19:        end if
20:     end for
21:     while $T_{p}$ is not empty do
22:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
23:        if $rand<u$ then
24:           $r\leftarrow T_{p}$.get_next_request()
25:        else
26:           $r\leftarrow T_{f}$.get_next_request()
27:        end if
28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
29:        if $l>0$ then
30:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
31:           if $rand<u$ then
32:              $T_{f}$.remove_request($r$)
33:           else
34:              $T_{p}$.remove_request($r$)
35:           end if
36:        else
37:           break
38:        end if
39:     end while
40:     return $B$
41:  end function

### v2

@@
 Abstract
-Recent advancements in large language models (LLMs) have facilitated a wide range of applications with distinct quality-of-experience requirements, from latency-sensitive online tasks, such as interactive chatbots, to throughput-focused offline tasks like document summarization.
-While deploying dedicated machines for these services ensures high-quality performance, it often results in resource underutilization.
+Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like document summarization.
+The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
 This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements.
-HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor for batch execution time estimation and an SLO-aware profiler to quantify interference, and (2) SLO-aware offline scheduling policies that maximize throughput and prevent starvation, without compromising online serving latency.
-Our evaluation on production workloads shows that HyGen achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency.
-The code will be released soon.
+HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor to estimate batch execution time and an SLO-aware profiler to quantify latency interference, and (2) SLO-aware offline scheduling policies that maximize serving throughput and prevent starvation, without compromising online serving latency.
+Our evaluation on production workloads shows that HyGen achieves up to 3.87$\times$ overall throughput and 5.84$\times$ offline throughput gains over online and hybrid serving baselines, respectively, while strictly satisfying latency SLOs.
 1 Introduction
-Large language models (LLMs) have emerged as transformative tools across diverse domains, powering applications from interactive chatbots to document analysis systems Grand View Research (2023).
-As LLM services continue to spread, their serving deployment must handle both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al.
+Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al.
 (2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)).
-Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), to ensure smooth user interactions.
-In contrast, offline tasks prioritize high throughput and resource utilization, often processing large batches of requests with relaxed latency constraints.
+Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints.
 The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al.
@@
 (2024).
-However, real-world LLM workloads demonstrate significant temporal variations in request loads.
+However, real-world LLM workloads demonstrate significant temporal variations in request load.
 Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3).
-To maintain acceptable latency under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
+To meet latency requirements under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
 This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster.
-By opportunistically co-locating offline tasks with online tasks during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests.
+By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests.
 However, realizing this opportunity requires addressing several fundamental challenges.
@@
 Second, LLM workloads are inherently unpredictable in both their arrival patterns and resource demands.
-Online requests arrive in unexpected bursts, while offline tasks appear in irregular batch intervals.
+Request arrival rates exhibit both diurnal patterns and unpredictable short-term fluctuations.
 This variability is further complicated by uncertainty in resource demands—input sequences vary widely in length, and the number of output tokens can hardly be predicted until generation completes.
-Third, co-locating online and offline workloads introduces interference, which degrades the performance of both.
-For instance, offline requests can delay time-sensitive online requests, while using large batch sizes to boost offline request throughput can negatively impact the latency of online requests.
+Third, co-locating online and offline workloads introduces interference.
+For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve offline request throughput can increase the latency of online requests.
 Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
 This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads.
-HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of heterogeneous request batches, (2) an interference-aware profiler that quantifies the performance impact of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online serving.
-Our evaluation on production workloads shows that HyGen guarantees latency requirements in co-locating requests and achieves up to 5.84$\times$ higher throughput compared to existing advances while maintaining comparable latency.
+HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of different request batches, (2) an interference-aware profiler that quantifies the performance interference of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online requests.
+Our evaluation on production workloads shows that HyGen improves serving throughput by 3.87-5.84$\times$ over existing advances Agrawal et al.
+(2024), while guaranteeing strict SLO compliance.
 To summarize, this paper makes the following contributions:
   • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
-  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with guaranteed performance isolation.
-  • To improve the performance isolation and throughput of HyGen, we design performance control mechanisms and SLO-aware offline scheduling policies.
+  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with performance isolation.
   • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
@@
 (2023); Liu et al.
-(2023), and model compression through quantization approaches Xiao et al.
+(2023), and model compression approaches Xiao et al.
 (2023); Lin et al.
@@
 Orca Yu et al.
-(2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching, and enables early exits for completed requests, thereby reducing batching latency.
+(2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching and enables early exits for completed requests.
 vLLM Kwon et al.
-(2023) develops PagedAttention for efficient memory management, while Sarathi-Serve proposes chunked prefill to minimize inter-request interference within batches.
-However, few existing systems explicitly address the challenge of co-locating latency-critical online requests with offline workloads to maximize resource utilization—a key focus of our work.
+(2023) develops PagedAttention for efficient memory management, while Sarathi-Serve Agrawal et al.
+(2024) proposes chunked prefill to minimize inter-request interference within batches.
+However, few existing systems address the challenge of co-locating latency-critical online requests with offline workloads.
 Throughput-Oriented LLM Inference.
@@
 FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput.
-For KV cache optimization, InfiniGen Lee et al.
-(2024) proposes dynamic cache management through selective prefetching based on token importance prediction.
 S3 Jin et al.
-(2023) tackles the KV cache problem through output sequence length prediction, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
-NanoFlow Zhu et al.
-(2024) focuses on maximizing hardware utilization through intra-device parallelism, introducing nano-batching and execution unit scheduling to overlap compute-, memory-, and network-intensive operations.
+(2023) utilizes output sequence length prediction technique, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
 Co-locating Workloads.
-Co-locating latency-critical applications with batch applications is a widely adopted approach to improve resource utilization in datacenters, with the popular resource partitioning techniques Chen et al.
-(2019b); Zhu & Erez (2016); Lo et al.
-(2015); Wu & Martonosi (2008); Chen et al.
-(2019a); Kasture et al.
-(2015); Sanchez & Kozyrakis (2011); Kasture & Sanchez (2014); Zhang et al.
+Co-locating latency-sensitive applications with batch applications has been widely adopted to improve resource utilization in datacenters Chen et al.
+(2019); Zhu & Erez (2016); Zhang et al.
 (2019); Iorgulescu et al.
@@
 (2023).
-Prior work has also explored multiplexing to serve multiple DNNs Dhakal et al.
+Prior work has also explored resource multiplexing to serve multiple DNNs Dhakal et al.
 (2020); Tan et al.
@@
 3.1 LLM Serving
-LLM generation consists of two sequential stages: (i) Prefilling Stage: It processes the entire input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) of the input sequence to avoid regeneration in subsequent tokens, and then the (ii) Decoding Stage: the model generates tokens one at a time in an auto-regressive manner.
+LLM generation consists of two sequential stages: (1) Prefill Stage: It processes the input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) to avoid regeneration in subsequent tokens, and the (2) Decode Stage: the model generates tokens one at a time in an auto-regressive manner.
 Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead.
-This stage iterates until meeting a termination condition, such as generating an end-of-sequence token or reaching the maximum length.
-Model executions in the prefilling stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously.
-Yet in the decoding phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al.
+Model executions in the prefill stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously.
+Yet in the decode phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al.
 (2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
@@
 This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
-The second counterpart is offline serving, which prioritizes throughput over latency, measured in requests per second (RPS) or tokens per second (TPS) for prefill/decode operations.
-Key applications include model benchmarking Hendrycks et al.
+The second scenario is offline serving, which prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS).
+For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs OpenAI (2024).
+Applications include model benchmarking Hendrycks et al.
 (2021b); Meta-Team (2024), document processing Chen et al.
@@
 3.2 Motivation and Challenges
-(a) 1 hour
-(b) 2 minutes
-Figure 1: Request rate variation in Microsoft Azure’s LLM service over one-hour and two-minute periods, showing significant temporal fluctuations between peak and off-peak periods.
+(a) One-hour Request Rates
+(b) Two-minute Request Rates
+Figure 1: Request rate varies significantly in Microsoft Azure’s LLM service over one-hour and two-minute periods.
 LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads.
@@
 This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
-This load variability suggests an opportunity to improve GPU utilization by co-locating online inference requests with offline batch workloads.
+This load variability suggests an opportunity to improve GPU utilization by co-locating online requests with offline requests.
 During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources.
-However, co-locating these workloads introduces several fundamental challenges:
+While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location Agrawal et al.
+(2024), doing so at scale introduces several fundamental challenges:
   1. 1.
     Diverse Latency Requirements: Applications and even requests of an application have distinct latency requirements.
-For example, users of different ages have different reading speeds, thus TBT requirements Liu et al.
-(2024a).
-These requirements are often expressed through multiple statistical metrics - some services prioritize tail latency guarantees (e.g., P99 TTFT $\leq$ 1000ms), others focus on average performance (e.g., mean TBT $\leq$ 50ms), and many need to satisfy both simultaneously.
-How to provide latency guarantees for diverse requests while sharing resources?
+For example, paid users require strict latency SLOs while free users accept more relaxed guarantees.
+How to flexiblely specify latency guarantees for online requests when sharing resources?
   2. 2.
     Massive Uncertainties: LLM serving faces temporal uncertainty—online requests arrive in unpredictable bursts with varying urgency levels—as well as resource demand uncertainty due to unpredictable output lengths.
-A seemingly simple query might trigger an elaborate response requiring thousands of tokens.
 How to perform efficient scheduling in the wild?
   3. 3.
-    Request Interference: Co-locating online and offline workloads introduces interference that degrades system performance.
-Large batches of offline requests arriving simultaneously can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests.
-Worse, batching requests of long inputs (e.g., a long document summarization request with 2,048 tokens) with short interactive queries (e.g., 128 tokens) can result in a dramatic increase in attention computation overhead, elongating the latency for all requests in the batch by an order of magnitude Agrawal et al.
+    Request Interference: Co-locating online and offline workloads introduces performance interference.
+Large batches of offline request can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests.
+Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude Agrawal et al.
 (2024).
@@
 4.1 Overview
-Figure 2: HyGen overview: Online requests and offline requests sent by users are asynchronously selected and processed.
+Figure 2: HyGen overview. Online requests and offline requests are asynchronously processed. Offline requests are opportunistically scheduled while retaining the batch latency budget.
 In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees.
@@
 2:     Input: running requests $R$, request queue $Q$,
-3:      latency budget $t$, remaining chunk size $c$,
-4:      memory budget $m$
-5:     Output: batched requests $B$
-6:     $B\leftarrow\{\}$
-7:     for $r\in R$.decoding do
-8:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
-9:        if $t_{req}\leq t$ then
-10:           $t\leftarrow t-t_{req}$
-11:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
-12:        end if
-13:     end for
-14:     for $r\in R$.prefilling $\cup Q$ do
-15:        TRY_SCHEDULE:
-16:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-17:        if $l>0$ then
-18:           $t\leftarrow t-t_{req}$
-19:           $c\leftarrow c-l$
-20:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
-21:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-22:        else
-23:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
-24:              PERFORM_PREEMPTION($R$, $m$)
-25:              goto TRY_SCHEDULE
-26:           else
-27:              break
+3:      latency budget $t$, chunk size $c$, memory budget $m$
+4:     Output: batched requests $B$
+5:     $B\leftarrow\{\}$
+6:     for $r\in R$.decode do
+7:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
+8:        if $t_{req}\leq t$ or PHASE ==absent$=$ = ONLINE then
+9:           $t\leftarrow t-t_{req}$
+10:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
+11:        end if
+12:     end for
+13:     for $r\in R$.prefill $\cup Q$ do
+14:        TRY_SCHEDULE:
+15:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+16:        if $l>0$ then
+17:           $t\leftarrow t-t_{req}$
+18:           $c\leftarrow c-l$
+19:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
+20:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
+21:        else
+22:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
+23:              PERFORM_PREEMPTION($R$, $m$)
+24:              goto TRY_SCHEDULE
+25:           else
+26:              break
+27:           end if
 28:        end if
-29:        end if
-30:     end for
-31:     return $B$
-32:  end function
-Figure 2 illustrates the system architecture of HyGen.
-To handle diverse latency requirements and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector that tracks active requests and their states.
-This separation allows HyGen to support scheduling policies for different request types without the complexity of managing mixed priorities in a single queue.
-The scheduling process in HyGen operates in two consecutive phases to carefully manage interference between online and offline requests.
-In the online phase, the scheduler processes the online request queue to construct an initial batch based on latency SLOs.
-This phase follows traditional continuous batching principles and supports various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al.
+29:     end for
+30:     return $B$
+31:  end function
+Figure 2 illustrates the overall design of HyGen.
+To handle diverse SLOs and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector.
+This design makes HyGen compatible with various existing scheduling policies for different request types.
+Following Sarathi-Serve Agrawal et al.
+(2024), we employ stall-free batching and chunked-prefills, batching prefill and decode requests together for efficient utilization of compute resources and dynamic co-location of requests.
+HyGen’s scheduling process occurs in two phases.
+The first (online) phase handles the online request queue to construct an initial batch, supporting various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al.
 (2023) and Virtual Token Counter (VTC) Sheng et al.
 (2024b).
-The scheduler prioritizes running requests before processing the waiting queue, continuing until reaching memory, batch size, or chunk size limits.
-A detailed workflow of this process is displayed in Algorithm 1.
-During a default two-phase scheduling, this function is called two times separately for online and offline requests.
-During online batch establishment, when the current online batch reaches the memory limit and more online requests remain to be scheduled, HyGen performs a priority-based preemption to maximize SLOs.
-To handle dynamic request arrivals while maintaining service guarantees, we implement an online-first preemption mechanism, where online requests are prioritized over all offline requests to minimize interference.
-Offline requests are prioritized based on their original priority in the waiting queue and their memory consumption.
-Preempted offline requests are removed from the current running list and reinserted into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
-After establishing the online batch, the scheduler enters the offline phase where it opportunistically fills the remaining capacity with offline requests.
-By utilizing a latency predictor to estimate batch execution time and determine remaining latency budget, instead of relying on static chunk size or batch size limits, we achieve both higher accuracy in interference prediction and strong transferability across different hardware configurations and workload patterns.
-This enables higher throughput while maintaining SLO guarantees by eliminating the need for conservative scheduling.
-The selection of offline requests then considers this available batch execution latency budget derived from user-specified SLOs, opportunities for efficient execution through prefix sharing, and current batch composition and resource constraints.
-Within this phase, running requests are prioritized over waiting ones to maintain execution efficiency.
-HyGen addresses these complex scheduling decisions through two key components detailed in the following sections.
-Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation decisions.
-Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining performance isolation through careful workload co-location.
-A detailed algorithm for the entire HyGen scheduling workflow can be found in Appendix A.1.
+If the memory limit is reached and more online requests remain, a priority-based preemption mechanism is employed by prioritizing online requests over offline requests to minimize interference.
+Preempted offline requests are reclaimed into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
+In the second (offline) phase, the remaining capacity is filled with offline requests using a latency predictor to capture the latency impact of request filling.
+The predictor calculates the latency of the online batch and selects offline requests based on the available batch execution latency budget, i.e., the latency slack without violating the latency requirement of online requests.
+It predicts the latency cost for decode requests, and calculates a maximum chunked prefill length for prefill requests for the given latency budget.
+Offline requests are scheduled in order until reaching the resource capacity, based on our SLO-aware scheduling policy.
+A formalized algorithm of the SLO-aware scheduling process is displayed in Algorithm 1.
+Appendix A.1 further presents a holistic workflow of two-phase scheduling, including the management of online and offline request queues and implementation of our SLO-aware scheduling algorithm.
+Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation.
+Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining SLOs.
 4.2 Performance Control Mechanisms
-Efficient co-location of online and offline requests requires precise control over execution timing and resource allocation.
-The key challenge lies in accurately estimating the performance impact of scheduling decisions while ensuring end-to-end SLO compliance.
-This section presents two core mechanisms in HyGen that address this challenge: a latency predictor for estimating batch execution times and an SLO-aware profiler for translating these estimates into scheduling decisions.
+Co-locating online and offline requests requires precise control over execution timing and resource allocation.
+The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure end-to-end SLO compliance.
+This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
 Latency Predictor.
 The design of our latency predictor is guided by three key requirements.
-First, it must provide fast inference to support real-time scheduling decisions when evaluating different batch compositions.
-Second, it needs to be robust across varying workload patterns conditions to maintain reliable performance.
-Third, it should be adaptable to different hardware configurations to capture their unique performance characteristics.
-The execution time of an LLM serving batch is primarily influenced by two distinct processing stages with different computational patterns.
+First, it must provide fast inference to support real-time scheduling decisions.
+Second, it needs to be robust across varying workload patterns to maintain reliable performance.
+Third, it should be adaptable to different hardware configurations to accurately capture their unique performance characteristics.
+The execution time of an LLM serving batch is primarily determined by two distinct processing stages with different computational patterns.
 The prefill stage exhibits quadratic complexity due to attention computations, with latency growing quadratically with input sequence length.
-This stage’s total load depends on both the number of requests and their individual sequence lengths.
+The total load of this stage depends on both the number of requests and their individual sequence lengths.
 In contrast, the decode stage shows linear scaling, with computational requirements growing proportionally with the number of tokens.
-To capture these characteristics, we model batch execution time as:
+We can model the batch execution time below to capture these characteristics:
 $T_{batch}=f(S_{p},S_{d},S_{p}^{2},S_{d}^{2},N_{p},N_{d})$ (1)
@@
 $N_{p}$ and $N_{d}$ represent the number of prefill and decode requests in the batch, respectively.
-We implement this model using linear regression, chosen for its combination of efficiency and effectiveness.
+We employ linear regression as the prediction model because of its efficiency and effectiveness.
+Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes.
 The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
-The training process involves three steps:
-  1. 1.
-    Training Data Collection: We systematically profile target hardware across diverse batch compositions, varying the number of requests in different phases, sequence length distributions, and total batch sizes.
-  2. 2.
-    Feature Engineering: For each batch, we compute six key features ($S_{p}$, $S_{d}$, $S_{p}^{2}$, $S_{d}^{2}$, $N_{p}$, $N_{d}$) based on token counts and batch sizes, capturing both linear and quadratic scaling.
-  3. 3.
-    Model Training: Linear regression coefficients are optimized to minimize prediction error across the training set, providing sufficient accuracy for scheduling decisions while maintaining fast inference times.
 SLO-Aware Profiling.
-While accurate latency prediction is essential, translating these predictions into scheduling decisions requires careful consideration of end-to-end performance requirements.
-Our SLO-aware profiler addresses this by determining appropriate latency budgets that ensure SLO compliance under various operating conditions.
-The profiler first analyzes specific workload-SLO combinations to establish viable latency budget ranges.
-Then, given that SLO has a positive correlation with batch latency, we perform the binary search on the batch latency budget to get an upper limit that meets the SLO.
+The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
+The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges.
+Given that a larger batch size and longer inputs will increase batch latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests.
+During deployment, this latency budget is then used as the batch latency limit in the two-phase SLO-aware scheduling process (Section 4.1) to ensure SLO compliance.
 This profiling enables three key capabilities.
@@
 Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
-Together, these mechanisms enable HyGen to make precise scheduling decisions that maximize resource utilization while maintaining strict performance guarantees.
-The latency predictor provides the quantitative basis for evaluating scheduling options, while the SLO-aware profiler ensures these decisions translate into consistent end-to-end performance.
 4.3 SLO-Aware Offline Scheduling Policies
-Offline scheduling in HyGen differs from online scheduling due to the variability of resource availability.
-Resources allocated to offline requests depend on the residual capacity after scheduling online requests, which fluctuates with the intensity of the online workload.
-Consequently, offline scheduling policies must dynamically adapt to available resources and optimize their specific objectives under such constraints.
-Common objectives include maximizing throughput, supporting multi-priority workloads, and, optionally, ensuring fairness Sheng et al.
-(2024b).
-The prioritization of these objectives varies depending on the chosen policy.
-Prefix sharing Gim et al.
-(2024); Gao et al.
+After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness Sheng et al. (2024b).
+To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy.
+Prefix sharing is a widely-adopted technique for reusing the KV cache of shared input prefix between requests Gim et al.
 (2024); Zheng et al.
-(2024a) has emerged as a widely-adopted technique to significantly improve throughput and reduce latency in LLM serving systems.
-The benefits are particularly pronounced in offline LLM serving, where requests can be strategically ordered to maximize prefix cache reuse Zhao et al.
+(2024a); Zhao et al.
 (2024); Zheng et al.
 (2024b).
-This optimization is especially effective for offline workloads since they have relaxed latency constraints, allowing the system to accumulate and reorder requests to maximize sharing opportunities.
-To showcase the extensibility and flexibility of HyGen’s design, we next introduce the SLO-aware Prefix Sharing Maximization (PSM) strategy, a throughput-optimized offline scheduling policy that leverages prefix reuse to improve resource efficiency.
-Recognizing the potential starvation issue inherent in naive PSM, we extend it with a fairness-enhanced variant.
-This improved PSM algorithm incorporates mechanisms such as freshness-aware scheduling and a unified utility function to balance throughput and fairness, mitigating starvation without compromising efficiency.
-The modular architecture of HyGen enables seamless integration of advanced offline scheduling policies.
-In addition to PSM, the system supports other strategies, such as a default First-Come-First-Serve (FCFS) policy and a Deadline-First policy for basic scheduling needs.
-This flexibility allows HyGen to cater to a wide range of objectives, including multi-priority support and fairness, making it a robust and adaptable solution for hybrid serving scenarios with dynamic and heterogeneous workloads.
+Figure 3: HyGen respects latency requirements in co-locating requests.
 Prefix Sharing Maximization Strategy.
-To fully leverage prefix sharing among offline requests, HyGen employs a Prefix Sharing Maximization (PSM) strategy designed for SLO-aware offline scheduling.
-This strategy organizes offline requests into a prefix tree and performs Depth-First Search (DFS) to generate a sorted waiting list.
-HyGen then performs SLO-aware scheduling on offline requests using this sorted list in a similar manner as introduced in Section 4.1, while maintaining the original DFS order to maximize prefix sharing.
-This enables reducing redundant computation and improving throughput.
-Despite its efficiency, the naive PSM strategy can lead to starvation for requests with minimal prefix-sharing potential.
-This issue can be mitigated by combining requests selected by PSM with those chosen based on freshness.
-The trade-off between these two selection methods can be managed using a utility ratio, ensuring a balance between efficiency and fairness.
-During scheduling, offline waiting requests are stored in both a prefix tree for maximum prefix sharing, and a self-balanced binary search tree (BST) sorted by freshness.
-When a new offline request can be scheduled from the waiting list, an offline request is selected either from the head of the queue in the DFS order of the prefix tree, or from the most stale request in the self-balanced BST, based on the utility value.
-The other data structure then syncs up by deleting the selected request from their lists, namely by prefix and index lookup.
-The formalized algorithm is detailed in Appendix A.2.
-While starvation due to the strict prioritization of online requests over offline ones is unavoidable to ensure performance isolation, the enhanced PSM algorithm effectively mitigates starvation caused by prefix-sharing optimizations.
-These enhancements make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
+Our Prefix Sharing Maximization (PSM) strategy organizes offline requests into a prefix tree that follows the structure of a Trie tree, capturing prefix sharing characteristics of all offline requests, with each leaf node representing an offline request.
+The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together, maximizing prefix cache reuse among requests.
+Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix sharing, reducing redundant computation and improving throughput.
+Additionally, the prefix tree structure ensures fast insertion of new arrivals and deletion of scheduled requests, guaranteeing efficient scheduling during runtime.
+A formalized algorithm is demonstrated in Appendix A.2.
+For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug).
+Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities.
+In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV-cache reuse through shared prefixes.
+The PSM strategy demonstrates strong extendability.
+Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential, blocked by continuous arrivals for an extended period.
+To illustrate this challenge, consider a request queue: (What is ML, What is AI, How to code, What is DL).
+When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely.
+This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness.
+Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness.
+A detailed algorithm is in Appendix A.3.
+These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
 5 Performance Evaluation
-Figure 3: HyGen respects latency requirements in co-locating requests.
-Figure 4: HyGen improves serving throughput under the same SLOs.
+Figure 4: HyGen improves serving throughput under varying SLOs.
 5.1 Evaluation Setup
-Implementation.
-We implement HyGen on top of the open-source implementation of vLLM Kwon et al.
+Implementation and Testbeds.
+We implement HyGen on top of vLLM Kwon et al.
 (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al.
 (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code.
-Testbeds.
-We evaluate HyGen on two server configurations.
-The first server is equipped with 4 NVIDIA A100 GPUs (40GB VRAM each), while the second server contains 4 NVIDIA A40 GPUs (48GB VRAM each).
-Both servers are identically configured with an AMD EPYC Milan CPU (1 socket, 64 cores, SMT disabled), 256GB DDR4 RAM, and a 1.5TB NVMe SSD (PCIe 4.0).
-Models.
-We used Llama2-7B Touvron et al.
-(2023) and Qwen-14B Bai et al.
-(2023) models to test the performance of HyGen on varying model structures and sizes.
-For the two models, we used A100 and A40 GPUs, respectively.
-Workloads.
+We evaluate HyGen on two server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each) and the other with 4 NVIDIA A40 GPUs (48GB VRAM each).
+Both servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
+Models and Workloads.
+We evaluated HyGen using Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models on A100 and A40 GPUs, respectively.
 For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al.
-(2024), a one-hour trace taken from the LLM services in Azure with real-world requests and timestamps.
-The request rate variation is shown in Figure 1.
-We randomly sampled the original trace to get the desired overall QPS that suits our hardware serving capacity.
-Specifically, for a time duration of $T$ seconds, we would sample $T\times Q$ requests from this time window to suit a desired QPS $Q$.
+(2024), a one-hour production trace with real-world requests and timestamps.
+We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity.
+Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$.
 For the offline workload, we used arXiv summarization Cohan et al.
-(2018), a dataset for summarization of long documents.
+(2018), a dataset for long document summarization.
 Metrics.
 For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values.
-For throughput evaluation, we evaluated our system on the maximum throughput of offline requests while meeting interference SLOs.
+We measure throughput using tokens per second (TPS) and queries per second (QPS).
 Baselines.
@@
 (2024) as our baseline.
-For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first two-phase scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.
-(2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to be an SLO-aware serving system, HyGen*.
+For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1.
+(2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*.
 Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference.
 The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
-5.2 End-to-end Performance
-HyGen respects latency requirements in co-locating requests.
-We tested HyGen on four SLO metrics, namely mean TBT, P99 TBT, mean TTFT, and mean TBT.
-For each metric, we set varying interference tolerance ratios as the SLO.
-Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings.
-Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
-Figure 5: Prefix sharing maximum scheduler achieves higher serving capacity.
-Figure 6: SLO-aware profiler bridges the gap between batch latency and overall QoE.
 Figure 5: HyGen latency predictor achieves high accuracy for batch latency prediction.
-Figure 8: HyGen meets stringent SLO under various QPS settings
+Figure 6: Prefix Sharing Maximization policy achieves higher serving capacity.
+Figure 7: SLO-aware profiler bridges the gap between batch latency and overall SLO.
 Figure 8: Hygen dynamically controls offline throughput according to online workload and latency budget.
@@
 Figure 9: Hygen meets SLOs and achieves higher throughput for Yi-34B model using TP=2, PP=2.
+Figure 10: HyGen meets SLO under various online QPS settings.
+Figure 11: HyGen is able to meet multiple SLOs simultaneously.
 (a) SLOs
@@
 (b) Throughput
-Figure 12: Hygen meets SLOs and achieves higher throughput for CNN/DM offline trace.
-Figure 13: HyGen is able to meet multiple SLOs.
-Figure 14: HyGen design is robust to predictor accuracy.
-Figure 15: Hygen adjusts offline throughput based on online QPS.
+Figure 13: Hygen meets SLOs and achieves higher throughput.
+Figure 14: HyGen is robust to predictor accuracy.
+Figure 15: Hygen dynamically adjusts offline throughput.
+5.2 End-to-end Performance
+HyGen respects latency requirements in co-locating requests.
+We tested HyGen on four SLO metrics, mean TBT, P99 TBT, mean TTFT, and P99 TTFT.
+For each metric, we set varying interference tolerance ratios as the SLO.
+Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings.
+Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
 HyGen improves serving throughput.
-Figure 4 showcases the offline throughput of HyGen for various metrics and tolerance ratios.
-Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.37×5.37$\times$5.37 × to 5.84×5.84$\times$5.84 × throughput gain.
-This verifies that our fine-grained latency predictor and SLO-aware profiler achieve higher offline serving efficiency compared to their simplified counterparts.
+Figure 4 shows the offline throughput of HyGen for various metrics and tolerance ratios.
+Through efficient request co-location, HyGen improves overall serving throughput by up to 3.87$\times$ compared to pure online serving.
+Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.84×5.84$\times$5.84 × offline throughput gain.
+This verifies that our fine-grained latency predictor and SLO-aware profiler designs achieve higher serving efficiency compared to their simplified counterparts.
 5.3 Performance Breakdown
@@
 We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset.
-Figure 7 shows the CDF of absolute error and absolute percentage error of the latency predictor.
-HyGen latency predictor is able to achieve a mean absolute percentage error of only 1.78% and 1.07%, verifying the high accuracy of our predictor design.
+Figure 5 shows that our latency predictor achieves a mean absolute percentage error of only 1.78% and 1.07%, confirming its high accuracy.
 Impact of prefix sharing.
-To test HyGen’s compatibility with prefix sharing, we did a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al.
-(2021a; b) as the offline dataset, running on Llama2-7B model.
-In our simulation, the shared prompt prefix length between two consecutively scheduled offline requests is deducted from the original prompt length of the latter one to simulate prefix sharing.
-We tested the offline throughput of HyGen with and without our prefix-aware offline schedule policy.
-Figure 6 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix-aware scheduling policy.
+To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al.
+(2021a; b) as the offline dataset on a Llama2-7B model.
+In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing.
+Figure 7 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix sharing maximization scheduling policy.
 Impact of SLO-aware profiler.
-To demonstrate the effect of the SLO-aware profiler in HyGen, we conducted experiments where we simply set the desired mean TBT SLO as the batch latency budget.
-Figure 6 shows the performance gap between individual batch latencies and overall mean TBT, and that our SLO-aware profiler is able to bridge this gap and provide hybrid serving with controlled SLO.
+To demonstrate the effect of HyGen’s SLO-aware profiler, we compared it with a simple strategy that sets the desired mean TBT SLO as the batch latency budget.
+Figure 7 shows the performance gap between individual batch latencies and overall mean TBT, illustrating how the SLO-aware profiler bridges this gap for controlled SLO in hybrid serving.
 Breakdown by time.
-Figure 9 shows a temporal breakdown of the throughput performance of HyGen during one serving process.
-At runtime, HyGen shows the ability to dynamically adjust offline throughput according to current online workload and overall latency budget.
-HyGen batches offline requests more aggressively during troughs of online QPS and lowers offline throughput during online bursts, harnessing compute resources in an adaptive manner.
+Figure 8 shows a temporal throughput breakdown of HyGen.
+At runtime, HyGen dynamically adjusts offline throughput based on online workload and overall latency budget, batching offline requests more aggressively during online QPS troughs and reducing offline throughput during online bursts, harnessing compute resources in an adaptive manner.
 5.4 Ablation Studies
 Impact of model parallelisms.
-To evaluate HyGen’s effectiveness in distributed inference scenarios, we deployed the Yi-34B model Young et al.
-(2024) using a hybrid parallelization strategy: tensor-parallelism (TP) Shoeybi et al.
+To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model Young et al.
+(2024) on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) Shoeybi et al.
 (2019) and pipeline-parallelism (PP) Athlur et al.
@@
 (2019) with degree 2 for each dimension.
-The online and offline workloads are Azure LLM Inference trace and arXiv summarization, respectively.
-Figure 10(a) and Figure 10 show that HyGen maintains its ability to meet SLOs and achieves higher throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
+Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
 Impact of SLO requirements.
-To further evaluate HyGen’s ability to meet stringent SLOs under various settings, we conducted experiments on the four aforementioned metrics under varied QPS settings, each with a 5% interference tolerance ratio.
-Figure 8 shows that HyGen is able to meet stringent SLOs for all metrics under various QPS settings.
+We further evaluate HyGen’s ability to meet stringent SLOs under varying online QPS settings on the four aforementioned metrics, each with 5% interference tolerance.
+Figure 10 shows that HyGen meets stringent SLOs for all metrics.
 We further demonstrate HyGen’s ability to meet multiple SLOs at the same time.
-Here, we tested HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%.
-Figure 13 shows that at a lower mean TBT tolerance, the performance of HyGen is bounded by mean TBT SLOs; however, after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control, demonstrating the flexibility of HyGen in meeting multiple SLOs.
+By testing HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%, Figure 11 shows that at a lower mean TBT tolerance, HyGen’s performance is bounded by mean TBT SLOs; after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control.
 Impact of models and datasets.
-We further evaluate HyGen on Mistral-7B and MooncakeQin et al.
-(2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace.
-The offline dataset used in the first experiment is arXiv summarization.
-For the second experiment, we replaced the offline dataset with CNN/DailyMail summarization dataset See et al.
+We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model with Mooncake trace Qin et al.
+(2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset See et al.
 (2017); Hermann et al.
-(2015) and used Llama2-7B and Azure LLM Inference trace as the model and online dataset, respectively.
-Figure 11 and Figure 12 show HyGen achieves superior performance than its counterparts under these two settings.
+(2015).
+An analysis of the Mooncake trace is shown in Appendix B.
+Figure 12 and Figure 13 show that HyGen achieves superior performance than its counterparts under these settings.
 Impact of predictor accuracy.
-To further evaluate the robustness of HyGen, we tested HyGen using several pre-trained LR latency predictors taken from other settings and workloads and tested them on Azure LLM Inference trace, resulting in varied prediction accuracy.
+We tested HyGen’s robustness using several pre-trained LR latency predictors with varying prediction accuracy taken from other workloads and tested them on Azure LLM Inference trace and arXiv summarization dataset.
 Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO.
-We notice that HyGen remains robust across different accuracy settings.
-Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on our testing CPU, guaranteeing transferability and negligible overhead.
-Impact of online arrival rates.
-Under high online request arrival rates, online requests may approach system serving capacity, leaving little headroom for offline request co-location.
-Figure 15 shows how online QPS affects offline throughput under a 5% P99 TBT SLO tolerance ratio.
-As online load increases, HyGen is able to adjust offline throughput according to the residual serving capacity of the system, while maintaining higher throughputs.
+HyGen remains robust across different accuracy settings.
+Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on CPUs.
+Impact of online arrival rate.
+Under a high online arrival rate, online requests may approach system serving capacity, limiting the headroom for request co-location.
+Figure 15 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance.
+As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs.
 6 Conclusion
 This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads.
-We employ a control mechanism to predict and manage interference impacts, while the scheduling framework supports multiple optimization objectives for offline serving.
-Evaluation on production workloads demonstrates that HyGen achieves up to 5.84$\times$ higher throughput while maintaining strict latency SLOs.
+We employ control mechanisms to predict and manage interference impacts, and a scheduling policy to opportunistically schedule offline serving.
+Evaluation on production workloads demonstrates that HyGen improves serving throughput by 3.87-5.84$\times$ while maintaining strict latency SLOs.
 Appendix
@@
 A.1 HyGen Two-phase Scheduling Algorithm
-This section gives a detailed explanation of the two-phase scheduling algorithm introduced in Section 4.
+This section gives a detailed and formalized demonstration of the two-phase scheduling algorithm introduced in Section 4.1.
+In each scheduling step, the workflow invokes the SLO-aware scheduling process in Algorithm 1 twice (line 13 and line 18) to form a hybrid batch with online and offline request co-location while respecting latency and memory limits.
 To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler.
@@
 The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget.
-The offline scheduling decisions are then sent back to the main process using the message queue and used for the coming scheduling step.
-To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization size $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
+The offline scheduling decisions are then sent back to the main process using the message queue and used for the next scheduling step.
+To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization degree $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
 Algorithm 2 HyGen two-phase scheduler
@@
 2:  global $R_{on}$ (online request running list), $R_{off}$ (offline request running list)
-3:  global $Q_{send}$, $Q_{recv}$ (message queue)
+3:  global $Q_{send}$, $Q_{recv}$ (message queues)
 4:  global $L$ (latency budget), $M$ (memory budget), $C$ (chunk size), $M_{off}$ (offline memory)
@@
 A.2 SLO-Aware Prefix Sharing Maximization Algorithm
-This section details the SLO-aware prefix sharing algorithm design in HyGen.
+This section details the SLO-aware prefix sharing maximization algorithm design in HyGen.
+We construct a prefix tree $T_{p}$ to capture prefix sharing charateristics among offline requests.
+During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled.
+Additionally, running requests keep their original DFS order in future scheduling process, effectively utilizing prefix sharing.
+Algorithm 3 Prefix-sharing-aware offline scheduler
+1:  Construct prefix tree $T_{p}$
+2:  function PREFIX_SHARING_OFFLINE_SCHEDULE
+3:     Input: running requests $R$, latency budget $t$, remaining chunk size $c$, memory budget $m$
+4:     Output: batched requests $B$
+5:     $B\leftarrow\{\}$
+6:     for $r\in R$ do
+7:        if $r$.state == DECODE then
+8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
+9:           if $t>t_{req}$ then
+10:              break
+11:           end if
+12:           $t\leftarrow t-t_{req}$
+13:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
+14:        else if $r$.state == PREFILL then
+15:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+16:           if $l>0$ then
+17:              $t\leftarrow t-t_{req}$
+18:              $c\leftarrow c-l$
+19:              $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
+20:              $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
+21:           else
+22:              break
+23:           end if
+24:        end if
+25:     end for
+26:     while $T_{p}$ is not empty do
+27:        $r\leftarrow T_{p}$.get_next_request()
+28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+29:        if $l>0$ then
+30:           $t\leftarrow t-t_{req}$
+31:           $c\leftarrow c-l$
+32:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
+33:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
+34:           $T_{p}$.remove_request($r$)
+35:        else
+36:           break
+37:        end if
+38:     end while
+39:     return $B$
+40:  end function
+A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
+This section details the extended version of the SLO-aware prefix sharing maximization algorithm, enhanced with fairness-aware scheduling.
 For offline requests, we construct a prefix tree $T_{p}$ for prefix sharing, and a self-balancing BST $T_{f}$ for request freshness.
-A utility value between 0 and 1 is used to balance the chance between these two data structures.
-During scheduling, prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, and keeps their original order when scheduling running requests, effectively utilizing prefix sharing while avoiding starvation.
+A utility ratio between 0 and 1 is used to balance the chance between these two data structures.
+During scheduling, the extended prefix sharing maximization algorithm retrieves waiting offline requests from either $T_{p}$ or $T_{f}$, based on the utility ratio.
+The selected request is then deleted from both data structures to ensure synchronization.
+This solution balances prefix sharing and fairness without disrupting the DFS order of the prefix tree while avoiding possible starvation.
 Algorithm 4 Prefix-sharing-aware offline scheduler
@@
 4:     Output: batched requests $B$
-5:     for $r\in R$ do
-6:        if $r$.state == DECODE then
-7:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
-8:           if $t>t_{req}$ then
-9:              break
-10:           end if
-11:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
-12:        else if $r$.state == PREFILL then
-13:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-14:           if $l>0$ then
-15:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-16:           else
-17:              break
-18:           end if
+5:     $B\leftarrow\{\}$
+6:     for $r\in R$ do
+7:        if $r$.state == DECODE then
+8:           $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
+9:           if $t>t_{req}$ then
+10:              break
+11:           end if
+12:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
+13:        else if $r$.state == PREFILL then
+14:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+15:           if $l>0$ then
+16:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
+17:           else
+18:              break
 19:           end if
-20:     end for
-21:     while $T_{p}$ is not empty do
-22:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
-23:        if $rand<u$ then
-24:           $r\leftarrow T_{p}$.get_next_request()
-25:        else
-26:           $r\leftarrow T_{f}$.get_next_request()
-27:        end if
-28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-29:        if $l>0$ then
-30:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-31:           if $rand<u$ then
-32:              $T_{f}$.remove_request($r$)
-33:           else
-34:              $T_{p}$.remove_request($r$)
-35:           end if
-36:        else
-37:           break
-38:        end if
-39:     end while
-40:     return $B$
-41:  end function
+20:        end if
+21:     end for
+22:     while $T_{p}$ is not empty do
+23:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
+24:        if $rand<u$ then
+25:           $r\leftarrow T_{p}$.get_next_request()
+26:        else
+27:           $r\leftarrow T_{f}$.get_next_request()
+28:        end if
+29:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+30:        if $l>0$ then
+31:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
+32:           $T_{p}$.remove_request($r$)
+33:           $T_{f}$.remove_request($r$)
+34:        else
+35:           break
+36:        end if
+37:     end while
+38:     return $B$
+39:  end function
+Appendix B Mooncake Trace Analysis
+This section provides an analysis of Mooncake trace, a production trace taken from real-world server with request lengths and request arrival timestamps.
+Figure 16 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods.
+This analysis further demonstrates the fluctuant characteristic of LLM services.
+(a) 1 hour
+(b) 10 minutes
+Figure 16: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.

### v4

@@
 Abstract
-Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like document summarization.
+Large language models (LLMs) have facilitated a wide range of applications with distinct service-level objectives (SLOs), from latency-sensitive online tasks like interactive chatbots to throughput-oriented offline workloads like data synthesis.
 The existing deployment model, which dedicates machines to each workload, simplifies SLO management but often leads to poor resource utilization.
-This paper introduces HyGen, an interference-aware LLM serving system that enables efficient co-location of online and offline workloads while preserving latency requirements.
-HyGen incorporates two key innovations: (1) performance control mechanisms, including a latency predictor to estimate batch execution time and an SLO-aware profiler to quantify latency interference, and (2) SLO-aware offline scheduling policies that maximize serving throughput and prevent starvation, without compromising online serving latency.
-Our evaluation on production workloads shows that HyGen achieves up to 3.87$\times$ overall throughput and 5.84$\times$ offline throughput gains over online and hybrid serving baselines, respectively, while strictly satisfying latency SLOs.
+[…]
 1 Introduction
-Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions OpenAI (2022); Liu et al.
-(2024a)) and throughput-oriented offline tasks (e.g., document summarization Grand View Research (2023)).
+Large language models (LLMs) have emerged as transformative tools across diverse domains, handling both latency-critical online requests (e.g., chatbot interactions andes ; chatgpt ) and throughput-oriented offline tasks (e.g., document summarization llm-market-report ).
 Online serving demands low and stable response times, measured by Time to First Token (TTFT) and Time Between Tokens (TBT), while offline tasks prioritize high throughput and resource utilization, often processing large batches with relaxed latency constraints.
-The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate GPU clusters Sheng et al.
-(2023); Zhao et al.
-(2024); Patel et al.
-(2024); Stojkovic et al.
-(2024).
-However, real-world LLM workloads demonstrate significant temporal variations in request load.
-Our analysis of production traces shows that in addition to the popular diurnal request arrival patterns, online request rates can vary by up to 3$\times$ within minutes (Section 3).
-To meet latency requirements under such bursty loads, service providers must provision GPU resources for peak demand, leading to substantial underutilization during off-peak hours.
-This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same GPU cluster.
+The disparity in these requirements has led most production deployments to segregate online and offline serving onto separate clusters to avoid interference patel2024splitwise ; sheng2023flexgen ; stojkovic2024dynamollm ; zhao2024blendserve .
+[…]
+This resource overprovisioning suggests an opportunity to improve resource efficiency via hybrid serving—co-locating online and offline workloads on the same inference engine instance.
 By opportunistically padding online requests with offline requests during periods of low online load, the system could maintain high GPU utilization while preserving latency guarantees for online requests.
 However, realizing this opportunity requires addressing several fundamental challenges.
-First, LLM services exhibit diverse latency requirements across applications Liu et al.
-(2024a).
+First, LLM services exhibit diverse latency requirements across applications andes .
 Interactive chatbots require consistent response times with both low initial latency (TTFT) and smooth token generation (TBT), while batch processing tasks prioritize throughput over latency.
@@
 Third, co-locating online and offline workloads introduces interference.
-For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve offline request throughput can increase the latency of online requests.
-Effectively managing this interference while ensuring performance guarantees demands meticulous orchestration of resource sharing.
+For instance, offline requests may delay time-sensitive online requests; using large batch sizes to improve the throughput of offline requests can increase the latency of online requests.
+Effectively managing this interference while ensuring latency guarantees demands meticulous orchestration of resource sharing.
 This paper presents HyGen, an interference-aware LLM serving system that elastically co-locates online and offline workloads.
 HyGen introduces several key techniques: (1) a latency predictor that accurately estimates the execution time of different request batches, (2) an interference-aware profiler that quantifies the performance interference of co-location, and (3) an adaptive scheduler that maximizes offline throughput while maintaining strict latency guarantees for online requests.
-Our evaluation on production workloads shows that HyGen improves serving throughput by 3.87-5.84$\times$ over existing advances Agrawal et al.
-(2024), while guaranteeing strict SLO compliance.
-To summarize, this paper makes the following contributions:
-  • We investigate the online request patterns in production LLM serving trace, revealing key challenges and opportunities in co-locating online and offline workloads.
-  • We propose HyGen, a new serving system that enables efficient resource sharing, improving throughput with performance isolation.
-  • We evaluate HyGen using production LLM serving traces to demonstrate HyGen’s effectiveness.
+[…]
+  • Key insights on the feasibility and benefits of co-locating online and offline LLM workloads, derived from systematic characterization of production traces.
+  • A statistical latency prediction model that accurately captures the relationship between batch composition and execution latency, accounting for quadratic complexity in prefill and linear scaling in decode phases.
+  • A novel scheduling system that dynamically co-locates online and offline workloads, formulated as a constrained optimization problem that maximizes throughput with prefix sharing while preserving strict latency SLOs and providing theoretical guarantees for fairness.
+[…]
 2 Related Work
-LLM Online Serving.
-Recent works have improved GPU computation through optimized kernels Dao et al.
-(2022); Ye et al.
-(2025), memory efficiency through efficient KV-cache management techniques Lin et al.
-(2024a); Liu et al.
-(2024b); Qin et al.
-(2024), distributed inference through various model parallelism strategies Shoeybi et al.
-(2019); Pope et al.
-(2023); Liu et al.
-(2023), and model compression approaches Xiao et al.
-(2023); Lin et al.
-(2024b); Dettmers et al.
-(2022).
-Orca Yu et al.
-(2022) introduces iteration-level scheduling to eliminate padding overhead in variable-length sequence batching and enables early exits for completed requests.
-vLLM Kwon et al.
-(2023) develops PagedAttention for efficient memory management, while Sarathi-Serve Agrawal et al.
-(2024) proposes chunked prefill to minimize inter-request interference within batches.
-However, few existing systems address the challenge of co-locating latency-critical online requests with offline workloads.
-Throughput-Oriented LLM Inference.
-As offline inference gains increasing traction, Huggingface Accelerate Hugging Face (2025), DeepSpeed ZeRO-Inference Aminabadi et al.
-(2022) and FlexGen Sheng et al.
-(2023) enable efficient LLM serving on commodity GPUs by offloading strategies.
-FlexGen aggregates memory from GPU, CPU and disk to maximize serving throughput.
-S3 Jin et al.
-(2023) utilizes output sequence length prediction technique, allowing more efficient memory allocation by avoiding over-allocation of the maximum sequence length.
-Co-locating Workloads.
-Co-locating latency-sensitive applications with batch applications has been widely adopted to improve resource utilization in datacenters Chen et al.
-(2019); Zhu & Erez (2016); Zhang et al.
-(2019); Iorgulescu et al.
-(2018); Patel & Tiwari (2020); Chen et al.
-(2023).
-Prior work has also explored resource multiplexing to serve multiple DNNs Dhakal et al.
-(2020); Tan et al.
-(2021); Zhang et al.
-(2023b), LLMs Duan et al.
-(2024), or LoRAs Chen et al.
-(2024); Sheng et al.
-(2024a); Wu et al.
-(2024).
-However, prior efforts have largely overlooked the opportunity to co-locate online and offline inference requests within the same LLM.
+LLM Inference Optimization.
+Recent advances in LLM inference through kernel optimization flashattention ; flashinfer , compilation frameworks ma2025intelligen ; wu2025mirage ; zheng2023einnet , and scheduling algorithms lin2024infinite ; lin2024parrot ; sun-etal-2025-disco have substantially improved LLM serving performance.
+Among these, a particularly important development is predictive scheduling, which aims to estimate request difficulty ding2024hybrid ; ong2024routellm or generation length fu2024efficient ; hu2024inference ; jin2023s ; patke2024queue ; qiu2024power .
+Contrasting these efforts, HyGen uniquely predicts batch execution time, providing precise control over request interference during co-located workloads.
+Additionally, the concept of request prefix sharing has been explored by gim2024prompt ; juravsky2024hydragen ; zheng2024sglang , where shared prefixes across requests optimize resource use.
+HyGen takes this a step further by applying a prefix sharing maximization strategy to opportunistically schedule offline requests, utilizing residual capacity from a primary, SLO-bound online workload.
+Moreover, HyGen introduces a fairness-aware extension to the standard prefix-sharing maximization (PSM) method, addressing the issue of starvation often encountered in naive designs.
+In parallel, the increasing demand for offline LLM inference has driven two key trends.
+First, platforms like Huggingface Accelerate hugging_face_accelerate , DeepSpeed ZeRO-Inference aminabadi2022deepspeed , and FlexGen sheng2023flexgen enable efficient inference on commodity hardware through memory offloading across GPUs, CPUs, and disk.
+Second, cloud providers, including OpenAI’s Batch API openai2024batchapi , have launched specialized services optimized for throughput rather than latency, processing millions of requests daily for large-scale data processing tasks.
+Workloads Co-location.
+In data center environments, co-locating latency-sensitive applications with batch applications has been explored to improve resource utilization chen2023olpart ; chen2019parties ; patel2020clite ; tempo-arixv25 .
+However, these approaches generally overlook the unique characteristics of LLM inference.
+In the context of LLM inference, several works have investigated the co-location of various model types and tasks.
+For instance, Punica chen2024punica , S-LORA sheng2024slora , and dLoRA wu2024dlora batch requests from different LoRAs, while MuxServe duanmuxserve multiplexes resources across multiple LLMs.
+In contrast, HyGen focuses on batching online and offline requests, addressing distinct optimization problems that are not interchangeable.
 3 Background and Motivation
-This section first introduces the core concepts of LLM serving systems (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
+This section first introduces the background of LLM serving deployment (Section 3.1), then illustrates how these characteristics motivate our system designs (Section 3.2).
 3.1 LLM Serving
-LLM generation consists of two sequential stages: (1) Prefill Stage: It processes the input prompt in parallel to generate the first output token, computing and caching intermediate model representations (i.e., the KV cache) to avoid regeneration in subsequent tokens, and the (2) Decode Stage: the model generates tokens one at a time in an auto-regressive manner.
-Each iteration processes only the most recent token while utilizing the KV cache of previous tokens, reducing computational overhead.
-Model executions in the prefill stage are compute-intensive due to the high degree of parallelism in processing the entire input sequence simultaneously.
-Yet in the decode phase, the same operations become memory-intensive due to the reliance on the KV cache Agrawal et al.
-(2024), and the total memory access cost grows linearly with the sequence length as each new token’s KV cache is appended.
-Serving Workloads.
-LLM serving deployment can be categorized into two primary scenarios.
-The first scenario is online serving, which targets real-time user interactions, such as chatbots, code assistants, and interactive applications OpenAI (2022); Microsoft (2023).
-This interactive nature often requires short Time to First Token (TTFT)—the latency from receiving a query to generating the first token—as well as the short Time Between Tokens (TBT), i.e., the inter-token generation latency in the decode stage.
-The second scenario is offline serving, which prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS).
-For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs OpenAI (2024).
-Applications include model benchmarking Hendrycks et al.
-(2021b); Meta-Team (2024), document processing Chen et al.
-(2021); Zhang et al.
-(2023a), data cleaning Narayan et al.
-(2022); Zheng et al.
-(2023), and synthetic data generation Allal et al.
-(2024).
+Large-scale inference clusters consist of multiple serving instances, with a router intelligently directing incoming requests to the most suitable instances srivatsa2024preble ; iccache-sosp25 .
+Each instance, which could be based on architectures like vLLM kwon2023efficient or SGLang zheng2024sglang , typically employs iteration-level scheduling yu2022orca and chunked prefill sarathi .
+This setup enables decode requests to perform an additional decoding step, while prefill requests are limited to a fixed token budget.
+These two types of requests are processed together within a single iteration, optimizing resource utilization.
+LLM serving deployment can be categorized into online serving and offline serving scenarios.
+Online serving targets real-time user interactions, such as chatbots, code assistants, and interactive applications githubGitHubCopilot ; chatgpt .
+This interactive nature often requires a short Time to First Token (TTFT) as well as a short Time Between Tokens (TBT).
+Offline serving prioritizes throughput over latency, measured in queries per second (QPS) or tokens per second (TPS).
+For example, OpenAI’s Batch API processes requests with relaxed latency requirements (up to 24 hours) at significantly lower costs compared to standard APIs openai2024batchapi .
+Applications include model benchmarking hendryckstest2021 ; dubey2024llama3herdmodels , document processing chen2021spreadsheetcoder ; zhang-etal-2023-summit , data cleaning narayan2022foundationmodelswrangledata ; zheng2023judgingllmasajudgemtbenchchatbot , and data synthesis huggingfaceCosmopediaCreate .
 3.2 Motivation and Challenges
-(a) One-hour Request Rates
-(b) Two-minute Request Rates
 Figure 1: Request rate varies significantly in Microsoft Azure’s LLM service over one-hour and two-minute periods.
+Figure 2: HyGen Overview. Online and offline requests are processed asynchronously, with offline requests opportunistically scheduled to respect latency budgets.
 LLM serving systems face a critical resource utilization challenge due to the highly variable nature of their workloads.
-Figure 1 reports our analysis of Microsoft Azure’s LLM service Patel et al.
-(2024), showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns.
-This variability creates an inherent tension in resource provisioning: GPU clusters must be sized to handle peak loads, inevitably leading to resource underutilization during off-peak periods.
-This load variability suggests an opportunity to improve GPU utilization by co-locating online requests with offline requests.
-During periods of low online traffic, the system could opportunistically schedule offline tasks to harvest idle GPU resources.
-While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location Agrawal et al.
-(2024), doing so at scale introduces several fundamental challenges:
+Figure 2 reports our analysis of Microsoft Azure’s LLM service patel2024splitwise , showing that request rates can fluctuate dramatically—varying up to 3× within minutes while following broader diurnal patterns.
+This variability creates an inherent tension in resource provisioning: serving clusters must be sized to handle peak loads, leading to resource underutilization during off-peak periods.
+This load variability suggests an opportunity to improve resource utilization by co-locating online requests with offline requests.
+For each serving instance in the cluster, during periods of low online traffic, it can opportunistically schedule offline tasks to harvest idle resources.
+While recent methodologies in simultaneous batching of prefill and decode requests set the premise for dynamic request co-location sarathi , doing so at scale introduces several fundamental challenges:
   1. 1.
@@
 For example, paid users require strict latency SLOs while free users accept more relaxed guarantees.
-How to flexiblely specify latency guarantees for online requests when sharing resources?
+How to respect diverse latency requirements in flight?
   2. 2.
@@
     Request Interference: Co-locating online and offline workloads introduces performance interference.
-Large batches of offline request can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests.
-Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude Agrawal et al.
-(2024).
+Large batches of offline requests can cause severe head-of-line blocking, delaying the processing of time-sensitive online requests.
+Worse, batching requests of long inputs with short interactive queries can elongate the latency of all requests in the batch by an order of magnitude sarathi .
 How to account for interference in co-locating requests?
@@
 4.1 Overview
-Figure 2: HyGen overview. Online requests and offline requests are asynchronously processed. Offline requests are opportunistically scheduled while retaining the batch latency budget.
-In this paper, we present HyGen, a novel LLM serving system that enables efficient co-location of online and offline requests while maintaining latency guarantees.
-Unlike conventional LLM serving systems that segregate online and offline workloads onto separate GPU clusters, HyGen coordinates both types of requests through a two-phase scheduling approach, maximizing resource utilization during off-peak periods while preserving performance isolation.
+HyGen introduces a novel approach to integrate online and offline requests while maintaining strict latency guarantees.
+As shown in Figure 2, HyGen employs a dual-queue architecture that separates latency-sensitive and throughput-oriented requests.
+This design accommodates diverse SLOs and variable workloads while remaining compatible with existing scheduling policies within each queue.
+We note that HyGen functions as an instance-level scheduler, receiving requests from an upstream system-level router (e.g., Preble srivatsa2024preble ).
+As a result, both the request concurrency and scheduling overhead at each instance are inherently bounded.
+HyGen’s two-phase scheduling operates as follows (see Appendix A.1 for the asynchronous two-queue workflow and message passing details):
+  • The online phase prioritizes latency-sensitive requests, forming an initial batch using established policies such as First-Come-First-Serve (FCFS) kwon2023efficient or fairness request scheduling sheng2024fairness .
+We introduce a priority-based preemption mechanism that protects online request performance by selectively preempting offline requests.
+Currently, HyGen preserves execution state for preempted requests, while its architecture supports various preemption mechanisms—including state discarding, preservation, and swapping—as categorized by InferCept abhyankar2024infercept .
+  • The offline phase uses our latency predictor to allocate remaining capacity to throughput-oriented requests.
+This predictor accurately estimates the latency impact of each potential offline request addition, determining either decode request latency costs or maximum chunked prefill lengths that fit within the available latency budget without violating online SLOs.
+Algorithm 1 formalizes our scheduling approach, with implementation details and complexity analysis ($O(n)$ where $n$ is the number of requests) provided in Appendices A.1 and A.4.
+Appendix C presents a novel cluster serving paradigm that addresses the longstanding tradeoff between SLO compliance and resource utilization based on HyGen.
 Algorithm 1 HyGen SLO-aware scheduler
@@
 6:  for $r\in R$.decode do
-7:        $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)
-8:        if $t_{req}\leq t$ or PHASE ==absent$=$ = ONLINE then
-9:           $t\leftarrow t-t_{req}$
-10:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
-11:        end if
-12:     end for
-13:     for $r\in R$.prefill $\cup Q$ do
-14:        TRY_SCHEDULE:
-15:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+7:   $t_{req}\leftarrow$ PREDICTOR.predict($r$, DECODE)    // predict latency of the decoding request
+8:   if $t_{req}\leq t$ or PHASE ==== ONLINE then
+9:    // schedule request if it is: 1. online, or 2. offline and enough latency budget left
+10:    $t\leftarrow t-t_{req}$
+11:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
+12:  for $r\in R$.prefill $\cup Q$ do
+13:   TRY_SCHEDULE:    // try to schedule a prefilling or waiting request
+14:   // get the max number of tokens allowed under memory and latency budget
+15:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_tokens($t$, $c$, $m$, $r$)
 16:   if $l>0$ then
-17:           $t\leftarrow t-t_{req}$
-18:           $c\leftarrow c-l$
-19:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
-20:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-21:        else
-22:           if PHASE ==absent$=$ = ONLINE and $R\neq B$ then
-23:              PERFORM_PREEMPTION($R$, $m$)
-24:              goto TRY_SCHEDULE
-25:           else
-26:              break
-27:           end if
-28:        end if
-29:     end for
-30:     return $B$
-31:  end function
-Figure 2 illustrates the overall design of HyGen.
-To handle diverse SLOs and unpredictable workloads, HyGen maintains separate queues for online and offline requests, each managed by a dedicated request selector.
-This design makes HyGen compatible with various existing scheduling policies for different request types.
-Following Sarathi-Serve Agrawal et al.
-(2024), we employ stall-free batching and chunked-prefills, batching prefill and decode requests together for efficient utilization of compute resources and dynamic co-location of requests.
-HyGen’s scheduling process occurs in two phases.
-The first (online) phase handles the online request queue to construct an initial batch, supporting various scheduling policies such as First-Come-First-Serve (FCFS) Kwon et al.
-(2023) and Virtual Token Counter (VTC) Sheng et al.
-(2024b).
-If the memory limit is reached and more online requests remain, a priority-based preemption mechanism is employed by prioritizing online requests over offline requests to minimize interference.
-Preempted offline requests are reclaimed into the offline queue, where they maintain their execution state to enable efficient rescheduling in subsequent batches.
-In the second (offline) phase, the remaining capacity is filled with offline requests using a latency predictor to capture the latency impact of request filling.
-The predictor calculates the latency of the online batch and selects offline requests based on the available batch execution latency budget, i.e., the latency slack without violating the latency requirement of online requests.
-It predicts the latency cost for decode requests, and calculates a maximum chunked prefill length for prefill requests for the given latency budget.
-Offline requests are scheduled in order until reaching the resource capacity, based on our SLO-aware scheduling policy.
-A formalized algorithm of the SLO-aware scheduling process is displayed in Algorithm 1.
-Appendix A.1 further presents a holistic workflow of two-phase scheduling, including the management of online and offline request queues and implementation of our SLO-aware scheduling algorithm.
-Section 4.2 describes our performance control mechanisms, including the latency predictor and SLO-aware profiler that enable accurate resource allocation.
-Section 4.3 presents our SLO-aware offline scheduling policies that maximize throughput while maintaining SLOs.
+17:    // schedule request
+18:    $t\leftarrow t-t_{req}$
+19:    $c\leftarrow c-l$
+20:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
+21:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
+22:   else
+23:    if PHASE ==== ONLINE and $R\neq B$ then
+24:     PERFORM_PREEMPTION($R$, $m$)    // preempt request with lower priority
+25:     goto TRY_SCHEDULE    // try to schedule again
+26:    else
+27:     break
+28:  return $B$
 4.2 Performance Control Mechanisms
-Co-locating online and offline requests requires precise control over execution timing and resource allocation.
-The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure end-to-end SLO compliance.
+Co-locating online and offline requests under diverse SLO requirements requires precise control over resource allocation.
+The key challenge lies in accurately estimating the latency impact of scheduling decisions to ensure SLO compliance.
 This section introduces a latency predictor for estimating batch execution latency and an SLO-aware profiler for translating the estimates into scheduling decisions.
@@
 Training data for the model is collected by systematically profiling target hardware across diverse batch compositions, varying in the number of requests in different phases, sequence length distributions, and total batch sizes.
-The linear model enables rapid evaluation of different batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
+The linear model enables rapid evaluation of varying batch compositions during scheduling, while its simple feature set ensures stable predictions across varying conditions.
+Further discussion on the implementation and expandability of the LR predictor can be found in Appendix B.
 SLO-aware Profiling.
-The latency predictor provides accurate latency estimates for filling offline requests. Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
+The latency predictor provides accurate latency estimates for filling offline requests.
+Our SLO-aware profiler leverages a latency budget to ensure SLO compliance in scheduling.
 The profiler first analyzes the given combination of workload and SLO to establish viable latency budget ranges.
-Given that a larger batch size and longer inputs will increase batch latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests.
-During deployment, this latency budget is then used as the batch latency limit in the two-phase SLO-aware scheduling process (Section 4.1) to ensure SLO compliance.
-This profiling enables three key capabilities.
-First, it determines appropriate latency thresholds that maintain SLO compliance for different workload types.
-Second, it provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements.
-Third, it establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
+Given that larger batch sizes and longer inputs will increase latency, the profiler test-runs latency budgets within the range to check their compliance with the given SLO and employs binary search to decide an upper limit that meets the overall SLO for online requests.
+During deployment, this latency budget is used as the batch latency limit in the two-phase scheduling process (Section 4.1) to ensure SLO compliance.
+This profiling enables three key capabilities: (1) It determines appropriate latency thresholds that maintain SLO compliance for various workloads and limitations (e.g., power constraints patel2024splitwise ).
+(2) It provides flexible adaptation by adjusting budgets based on changing workload characteristics and performance requirements.
+(3) It establishes a robust foundation for hybrid scheduling by accounting for both online and offline workload patterns.
 4.3 SLO-aware Offline Scheduling Policies
-After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness Sheng et al. (2024b).
+After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness sheng2024fairness .
 To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy.
-Prefix sharing is a widely-adopted technique for reusing the KV cache of shared input prefix between requests Gim et al.
-(2024); Zheng et al.
-(2024a); Zhao et al.
-(2024); Zheng et al.
-(2024b).
-Figure 3: HyGen respects latency requirements in co-locating requests.
+Prefix sharing is a widely adopted technique for reusing the KV cache of shared input prefixes between requests gim2024prompt ; zhao2024blendserve ; zheng2024sglang ; zheng2024batchllm .
 Prefix Sharing Maximization Strategy.
-Our Prefix Sharing Maximization (PSM) strategy organizes offline requests into a prefix tree that follows the structure of a Trie tree, capturing prefix sharing characteristics of all offline requests, with each leaf node representing an offline request.
-The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together, maximizing prefix cache reuse among requests.
-Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix sharing, reducing redundant computation and improving throughput.
-Additionally, the prefix tree structure ensures fast insertion of new arrivals and deletion of scheduled requests, guaranteeing efficient scheduling during runtime.
-A formalized algorithm is demonstrated in Appendix A.2.
+Our PSM strategy organizes offline requests into a prefix tree following the structure of a Trie tree with each leaf node representing a request, capturing prefix sharing characteristics of all offline requests.
+The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together.
+Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix cache reuse, reducing redundant computation and improving throughput.
+The prefix tree structure also ensures fast insertion and deletion in runtime scheduling.
+A formalized algorithm is in Appendix A.2.
 For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug).
 Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities.
-In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV-cache reuse through shared prefixes.
+In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV cache reuse through shared prefixes.
 The PSM strategy demonstrates strong extendability.
-Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential, blocked by continuous arrivals for an extended period.
-To illustrate this challenge, consider a request queue: (What is ML, What is AI, How to code, What is DL).
+Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential.
+Consider a request queue: (What is ML, What is AI, How to code, What is DL).
 When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely.
@@
 5 Performance Evaluation
+Figure 3: HyGen respects latency requirements in co-locating requests.
 Figure 4: HyGen improves serving throughput under varying SLOs.
-5.1 Evaluation Setup
-Implementation and Testbeds.
-We implement HyGen on top of vLLM Kwon et al.
-(2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al.
-(2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code.
-We evaluate HyGen on two server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each) and the other with 4 NVIDIA A40 GPUs (48GB VRAM each).
-Both servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
-Models and Workloads.
-We evaluated HyGen using Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models on A100 and A40 GPUs, respectively.
-For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al.
-(2024), a one-hour production trace with real-world requests and timestamps.
-We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity.
-Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$.
-For the offline workload, we used arXiv summarization Cohan et al.
-(2018), a dataset for long document summarization.
-Metrics.
-For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values.
-We measure throughput using tokens per second (TPS) and queries per second (QPS).
-Baselines.
-For pure online inference, we use Sarathi Agrawal et al.
-(2024) as our baseline.
-For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1.
-(2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*.
-Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference.
-The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
 Figure 5: HyGen latency predictor achieves high accuracy for batch latency prediction.
-Figure 6: Prefix Sharing Maximization policy achieves higher serving capacity.
-Figure 7: SLO-aware profiler bridges the gap between batch latency and overall SLO.
-Figure 8: Hygen dynamically controls offline throughput according to online workload and latency budget.
+Figure 6: Prefix Sharing Maximization improves serving throughput.
+Figure 7: SLO-aware profiler contributes to HyGen’s improvements.
+Figure 8: HyGen dynamically controls throughput according to online workload.
 (a) SLOs
@@
 (b) Throughput
-Figure 12: Hygen meets SLOs and achieves higher throughput for Mooncake trace and Mistral-7B model.
+Figure 12: HyGen meets SLOs and achieves higher throughput with CNN/DailyMail offline dataset.
+(a) 1 hour
+(b) 10 minutes
+Figure 13: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.
 (a) SLOs
 (b) Throughput
-Figure 13: Hygen meets SLOs and achieves higher throughput.
+Figure 14: HyGen meets SLOs and achieves higher throughput for Mooncake trace.
+(a) SLOs
+(b) Throughput
+Figure 15: HyGen meets SLOs and achieves higher throughput on A5000 GPU and Sheared-LLaMA-2.7B model.
 Figure 16: HyGen is robust to predictor accuracy.
 Figure 17: HyGen dynamically adjusts offline throughput.
+5.1 Evaluation Setup
+Implementation and Testbeds.
+We implement HyGen on top of vLLM kwon2023efficient ; vllm2023github and Sarathi sarathi ; sarathi2024github , with 1,300 lines of additional code.
+We evaluate HyGen on three server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each), one with 4 NVIDIA A40 GPUs (48GB VRAM each), and one with 1 NVIDIA A5000 GPU (24GB VRAM).
+All servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
+Models and Workloads.
+For end-to-end evaluation, we use Llama2-7B touvron2023llama and Qwen-14B bai2023qwen models on A100 and A40 GPUs, respectively.
+Online workloads are based on the conversation trace from Azure LLM inference trace 2023 patel2024splitwise , a one-hour production trace with real-world requests and timestamps.
+We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity.
+Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$.
+For offline workloads, we use arXiv summarization cohan-etal-2018-discourse , a dataset for long document summarization.
+In our ablation studies, we evaluate HyGen across different model scales ranging from Sheared-LLaMA-2.7B xia2023sheared , Mistral-7B jiang2023mistral7b to Yi-34B young2024yi .
+The Mooncake trace mooncake is further used as the online trace, providing industrial request length distributions and arrival patterns, while CNN/DailyMail hermann2015teaching ; see-etal-2017-get and MMLU hendrycks2021ethics ; hendryckstest2021 are used as offline traces in the ablation studies.
+For interference evaluation, we focus on Time to First Token (TTFT) and Time Between Tokens (TBT), including their mean and 99th percentile (P99) values.
+Throughput is measured in tokens per second (TPS) and queries per second (QPS).
+Baselines.
+For pure online inference, we use Sarathi sarathi as our baseline.
+For pure offline serving, we use Sarathi-offline to evaluate the maximum offline serving capacity, where an optimal chunk size is profiled for offline workload to maximize throughput.
+The hyperparameter search of Sarathi-offline achieves $\sim$12% throughput gain compared to the default setup, ensuring optimal baseline performance for fair comparison.
+We then compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1.
+(2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*.
+Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference.
+The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
 5.2 End-to-end Performance
 HyGen respects latency requirements in co-locating requests.
-We tested HyGen on four SLO metrics, mean TBT, P99 TBT, mean TTFT, and P99 TTFT.
-For each metric, we set varying interference tolerance ratios as the SLO.
-Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings.
-Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
-HyGen improves serving throughput.
-Figure 4 shows the offline throughput of HyGen for various metrics and tolerance ratios.
-Through efficient request co-location, HyGen improves overall serving throughput by up to 3.87$\times$ compared to pure online serving.
-Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.84×5.84$\times$5.84 × offline throughput gain.
-This verifies that our fine-grained latency predictor and SLO-aware profiler designs achieve higher serving efficiency compared to their simplified counterparts.
+We evaluate HyGen under four SLO metrics (mean TBT, P99 TBT, mean TTFT, and P99 TTFT) with varied interference tolerance ratios.
+Figure 3 shows that HyGen controls interference and guarantees to meet specific SLOs across our settings.
+Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen shows efficient SLO-aware latency control.
+[…]
 5.3 Performance Breakdown
@@
 Impact of prefix sharing.
-To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al.
-(2021a; b) as the offline dataset on a Llama2-7B model.
+To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU hendrycks2021ethics ; hendryckstest2021 as the offline dataset on a Llama2-7B model.
 In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing.
-Figure 7 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix sharing maximization scheduling policy.
+Figure 7 shows that HyGen yields up to $4\times$ offline throughput gain with its prefix sharing maximization scheduling policy.
 Impact of SLO-aware profiler.
@@
 Impact of model parallelisms.
-To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model Young et al.
-(2024) on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) Shoeybi et al.
-(2019) and pipeline-parallelism (PP) Athlur et al.
-(2022); Narayanan et al.
-(2019); Huang et al.
-(2019) with degree 2 for each dimension.
-Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
+To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model young2024yi on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) shoeybi2019megatron and pipeline-parallelism (PP) athlur2022varuna ; huang2019gpipe ; narayanan2019pipedream with degree 2 for each dimension.
+Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to $1.89\times$) than the baseline.
 Impact of SLO requirements.
@@
 Impact of models and datasets.
-We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model with Mooncake trace Qin et al.
-(2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset See et al.
-(2017); Hermann et al.
-(2015).
-An analysis of the Mooncake trace is shown in Appendix B.
+We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model jiang2023mistral7b with Mooncake trace mooncake , a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset hermann2015teaching ; see-etal-2017-get .
+Figure 14 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods, further demonstrating the fluctuating nature of LLM services.
 Figure 12 and Figure 14 show that HyGen achieves superior performance than its counterparts under these settings.
+Impact of hardware testbeds.
+[…]
 Impact of predictor accuracy.
-We tested HyGen’s robustness using several pre-trained LR latency predictors with varying prediction accuracy taken from other workloads and tested them on Azure LLM Inference trace and arXiv summarization dataset.
-Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO.
-HyGen remains robust across different accuracy settings.
-Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on CPUs.
+[…]
 Impact of online arrival rate.
-Under a high online arrival rate, online requests may approach system serving capacity, limiting the headroom for request co-location.
 Figure 17 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance.
 As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs.
+Understandably, a high online arrival rate limits the headroom for co-location as it approaches system serving capacity.
 6 Conclusion
-This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads.
-We employ control mechanisms to predict and manage interference impacts, and a scheduling policy to opportunistically schedule offline serving.
-Evaluation on production workloads demonstrates that HyGen improves serving throughput by 3.87-5.84$\times$ while maintaining strict latency SLOs.
+[…]
+Acknowledgements
+We thank the anonymous reviewers for their constructive and insightful feedback.
+This work was supported in part by grants from Cisco and Google, and by an award from NVIDIA Academic Program.
+It also utilized the Delta system at the National Center for Supercomputing Applications (NCSA) through allocation CIS240236 from the ACCESS program.
 Appendix
@@
 15:    PREEMPT_OFFLINE($R_{off}$, $m$, $M_{off}$)
-16:        end if
-17:        $m\leftarrow m-M_{off}$
-18:        $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
-19:        $Q_{recv}$.send($B$, $t$, $m$)
-20:     end while
-21:  end function
-22:  function SCHEDULER
-23:     Output: batched requests $B$
-24:     scheduled requests $B$, latency budget $t$, memory budget $m$leftarrow $Q_{recv}$.get(block=True)
-25:     if REQUEST_ARRIVAL then
-26:        UPDATE_BUDGET($B$, $t$, $m$)
-27:        if $t<0$ or $m<0$ then
-28:           PREEMPT_UNTIL_FIT($B$, $t$, $m$)
-29:        end if
-30:     end if
-31:     $Q_{send}$.send($B$)
-32:     return $B$
-33:  end function
+16:   $m\leftarrow m-M_{off}$
+17:   $B\leftarrow B\cup$ SLO_AWARE_SCHEDULE($R_{off}$, $Q_{off}$, $t$, $c$, $m$)
+18:   $Q_{recv}$.send($B$, $t$, $m$)
+19:function SCHEDULER
+20:  Output: batched requests $B$
+21:  scheduled requests $B$, latency budget $t$, memory budget $m$ $\leftarrow$ $Q_{recv}$.get(block=True)
+22:  if REQUEST_ARRIVAL then
+23:   UPDATE_BUDGET($B$, $t$, $m$)
+24:   if $t<0$ or $m<0$ then
+25:    PREEMPT_UNTIL_FIT($B$, $t$, $m$)
+26:  $Q_{send}$.send($B$)
+27:  return $B$
 A.2 SLO-Aware Prefix Sharing Maximization Algorithm
 This section details the SLO-aware prefix sharing maximization algorithm design in HyGen.
-We construct a prefix tree $T_{p}$ to capture prefix sharing charateristics among offline requests.
+We construct a prefix tree $T_{p}$ to capture prefix sharing characteristics among offline requests.
 During scheduling, offline requests are selected in the DFS order of the prefix tree, and deleted once being scheduled.
@@
 10:     break
-11:           end if
-12:           $t\leftarrow t-t_{req}$
-13:           $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
-14:        else if $r$.state == PREFILL then
-15:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-16:           if $l>0$ then
-17:              $t\leftarrow t-t_{req}$
-18:              $c\leftarrow c-l$
-19:              $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
-20:              $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-21:           else
-22:              break
-23:           end if
-24:        end if
-25:     end for
-26:     while $T_{p}$ is not empty do
-27:        $r\leftarrow T_{p}$.get_next_request()
-28:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-29:        if $l>0$ then
-30:           $t\leftarrow t-t_{req}$
-31:           $c\leftarrow c-l$
-32:           $m\leftarrow m-$ GET_NUM_BLOCKS($l$)
-33:           $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-34:           $T_{p}$.remove_request($r$)
-35:        else
-36:           break
-37:        end if
-38:     end while
-39:     return $B$
-40:  end function
+11:    $t\leftarrow t-t_{req}$
+12:    $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
+13:   else if $r$.state == PREFILL then
+14:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+15:    if $l>0$ then
+16:     $t\leftarrow t-t_{req}$
+17:     $c\leftarrow c-l$
+18:     $m\leftarrow m$- GET_NUM_BLOCKS($l$)
+19:     $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
+20:    else
+21:     break
+22:  while $T_{p}$ is not empty do
+23:   $r\leftarrow T_{p}$.get_next_request()
+24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+25:   if $l>0$ then
+26:    $t\leftarrow t-t_{req}$
+27:    $c\leftarrow c-l$
+28:    $m\leftarrow m$- GET_NUM_BLOCKS($l$)
+29:    $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
+30:    $T_{p}$.remove_request($r$)
+31:   else
+32:    break
+33:  return $B$
 A.3 Extended SLO-Aware Prefix Sharing Maximization Algorithm
@@
 10:     break
-11:           end if
-12:           $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$($r$, 00, $t_{req}$)}
-13:        else if $r$.state == PREFILL then
-14:           $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-15:           if $l>0$ then
-16:              $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-17:           else
-18:              break
-19:           end if
-20:        end if
-21:     end for
-22:     while $T_{p}$ is not empty do
-23:        $rand\leftarrow$ RANDOM_NUMBER(0, 1)
-24:        if $rand<u$ then
-25:           $r\leftarrow T_{p}$.get_next_request()
-26:        else
-27:           $r\leftarrow T_{f}$.get_next_request()
-28:        end if
-29:        $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
-30:        if $l>0$ then
-31:           $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m-$ GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$($r$, $l$, $t_{req}$)}
-32:           $T_{p}$.remove_request($r$)
-33:           $T_{f}$.remove_request($r$)
-34:        else
-35:           break
-36:        end if
-37:     end while
-38:     return $B$
-39:  end function
-Appendix B Mooncake Trace Analysis
-This section provides an analysis of Mooncake trace, a production trace taken from real-world server with request lengths and request arrival timestamps.
-Figure 16 shows the varying request arrival rates of Mooncake trace over one-hour and ten-minute periods.
-This analysis further demonstrates the fluctuant characteristic of LLM services.
-(a) 1 hour
-(b) 10 minutes
-Figure 16: Request rate varies in Moonshot Mooncake’s LLM service over one-hour and ten-minute periods.
+11:    $t\leftarrow t-t_{req}$, $B\leftarrow B\cup$ {($r$, 0, $t_{req}$)}
+12:   else if $r$.state == PREFILL then
+13:    $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+14:    if $l>0$ then
+15:     $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
+16:    else
+17:     break
+18:  while $T_{p}$ is not empty do
+19:   $rand\leftarrow$ RANDOM_NUMBER(0, 1)
+20:   if $rand<u$ then
+21:    $r\leftarrow T_{p}$.get_next_request()
+22:   else
+23:    $r\leftarrow T_{f}$.get_next_request()
+24:   $l,t_{req}\leftarrow$ PREDICTOR.get_max_prefill($t$, $c$, $m$, $r$)
+25:   if $l>0$ then
+26:    $t\leftarrow t-t_{req}$, $c\leftarrow c-l$, $m\leftarrow m$- GET_NUM_BLOCKS($l$), $B\leftarrow B\cup$ {($r$, $l$, $t_{req}$)}
+27:    $T_{p}$.remove_request($r$)
+28:    $T_{f}$.remove_request($r$)
+29:   else
+30:    break
+31:  return $B$
+A.4 Complexity Analysis of the Two-Phase Scheduling Algorithm
+In this section, we present an analysis of the computational complexity of HyGen’s two-phase scheduler. The time complexity of its core components is as follows:
+  • Latency Prediction: $O(1)$ inference using a pre-trained LR model.
+  • PSM: The initial construction of the prefix tree is $O(N×L)$, where $N$ is the number of offline requests and $L$ is the average number of tokens.
+Each insertion or deletion costs $O(L)$.
+In implementation, getting the next request in the DFS order only takes $O(1)$, since the DFS order is put in a pre-processed list derived from the prefix tree and can be synced up with the prefix tree asynchronously.
+  • PSM with Fairness: In the fairness-aware PSM algorithm, a self-balancing BST is used for picking the stalest request.
+Each lookup, insertion, or deletion takes $O(\log n)$.
+In implementation, the requests can be kept in an FCFS queue, which syncs up with the BST asynchronously to guarantee correctness, so that each lookup still only takes $O(1)$ time.
+Overall, for the two-phase scheduling using existing policies, the time complexity remains $O(n)$ as asynchronous updates can be performed for advanced policies like fairness-aware PSM.
+Appendix B Further Discussion of the Latency Predictor
+Effective co-location in HyGen hinges on a latency predictor that is both highly accurate and computationally lightweight, as it enables real-time scheduling decisions without introducing significant overhead.
+To this end, we employ a linear regression (LR) model, which provides inference in constant time, $O(1)$, making it ideal for a dynamic serving environment.
+This section details the model’s formulation and its adaptability to the complexities of real-world deployment scenarios.
+Model Formulation.
+The execution time of a serving batch, $T_{\text{batch}}$, is primarily determined by the computational patterns of its prefill and decode stages.
+We model this relationship as a function of the batch’s composition:
+$T_{\text{batch}}=f(S_{p},S_{d},S_{p}^{2},N_{p},N_{d})$ (2)
+where $S_{p}$ and $S_{d}$ represent the total number of tokens in the prefill and decode phases, respectively, and $N_{p}$ and $N_{d}$ are the corresponding request counts.
+The quadratic term $S_{p}^{2}$ is crucial for capturing the non-linear scaling of the self-attention mechanism, which dominates the computational cost of the prefill stage.
+In contrast, the decode stage, which processes one token per request at a time, exhibits linear scaling with the batch size ($N_{d}$).
+The model’s coefficients are pre-trained on data gathered by systematically profiling the target hardware and LLM across a diverse set of batch compositions, a lightweight process that ensures applicability to any deployment environment.
+Robustness Through System-Level Design.
+While a linear model offers unparalleled efficiency, its accuracy in the face of dynamic system conditions (e.g., GPU temperature variations, resource contention) is a critical consideration.
+HyGen ensures robustness not through a more complex model, but through a synergistic system design.
+The SLO-aware profiler (Section 4.2) first establishes a macro-level operational latency budget, which implicitly captures the system’s current performance characteristics.
+The LR predictor then operates at a micro-level, making fine-grained decisions on batch composition within this pre-calibrated budget.
+This two-level approach effectively decouples the real-time scheduling decision from low-level hardware variability.
+As shown in Figure 17, this design allows the system to maintain robust performance and meet SLOs even in scenarios with predictor error rates exceeding 20%.
+Furthermore, the feature set is extensible; environmental factors like hardware load can be readily integrated into the model if required for specific use cases.
+Adaptability to Modern Model Architectures.
+The feature set defined in Equation 2 is sufficiently general to generalize across various modern LLM architectures without modification.
+For instance, for Mixture-of-Experts (MoE) models [48], where a fixed number of experts are activated per token, the resulting computational cost scales linearly with the total number of tokens processed and is effectively modeled by the $S_{p}$ and $S_{d}$ features.
+Similarly, in hybrid architectures that combine linear-complexity components (e.g., Mamba [19]) with quadratic-complexity attention, our model naturally captures both computational patterns; the linear cost is reflected in the learned coefficient for $S_{p}$, while the quadratic cost of the Transformer blocks is captured by the $S_{p}^{2}$ term.
+This adaptability underscores the predictor’s robust design, ensuring its relevance as LLM architectures continue to evolve.
+Appendix C Taming the Throughput-Latency Tradeoff in LLM Serving with HyGen
+Traditional instance scaling solutions address bursty workloads by launching new instances, which can cause tens of seconds to several minutes cold-start delays [15, 64, 66].
+To handle these delays, providers often keep standby instances online, leading to wasted resources during off-peak periods.
+In contrast, HyGen optimizes resource utilization by running offline workloads on idle resources and reallocating them to online requests in real-time, within a single inference iteration.
+This eliminates cold-start delays while ensuring high resource utilization.
+HyGen complements instance scaling solutions by automating the transition between online and offline workloads, reducing the need for manual intervention.
+While instance scaling manages large load fluctuations, HyGen ensures efficient resource use during low-traffic periods, optimizing overall system performance in fixed-size clusters.